### v1

# View-Invariant Pixelwise Anomaly Detection in Multi-object Scenes with Adaptive View Synthesis

Abstract
The inspection and monitoring of infrastructure assets typically requires identifying visual anomalies in scenes periodically photographed over time. Images collected manually or with robots such as unmanned aerial vehicles from the same scene at different instances in time are typically not perfectly aligned. Supervised segmentation methods can be applied to identify known problems, but unsupervised anomaly detection approaches are required when unknown anomalies occur. Current unsupervised pixel-level anomaly detection methods have mainly been developed for industrial settings where the camera position is known and constant. However, we find that these methods fail to generalize to the case when images are not perfectly aligned. We term the problem of unsupervised anomaly detection between two such imperfectly aligned sets of images as Scene Anomaly Detection (Scene AD). We present a novel network termed OmniAD to address the Scene AD problem posed. Specifically, we refine the anomaly detection method reverse distillation to achieve a 40% increase in pixel-level anomaly detection performance. The network’s performance is further demonstrated to improve with two new data augmentation strategies proposed that leverage novel view synthesis and camera localization to improve generalization. We validate our approach with qualitative and quantitative results on a new dataset, ToyCity, the first Scene AD dataset with multiple objects, as well as on the established single object-centric dataset, MAD. https://drags99.github.io/OmniAD/
Index Terms:
Anomaly Detection, Unsupervised Learning, Camera Localization
I Introduction
Figure 1: In Scene AD, we observe variations in camera positions between non-anomalous views (represented in green), and query views which may have anomalies in them (represented in red). In contrast, traditional AD methods typically feature aligned views between the non-anomalous views and anomalous views.
The inspection and monitoring of infrastructure assets typically requires identifying visual anomalies in scenes periodically photographed over time [1, 2, 3, 4, 5]. Images collected manually or with robots such as unmanned aerial vehicles from the same scene at different instances in time are typically not perfectly aligned. For instance, the United States Navy places significant emphasis on the monitoring and detection of anomalies during inspections of their ships, which is critical for maintaining security and operational readiness and safety [6, 7]. Similarly, bridge inspectors can leverage this technology to identify parts of the bridge that are changing or behaving abnormally. This information is vital for predicting future performance and for maintenance decision making. In each of these examples, new image sets are acquired at every inspection but typically unaligned with previous inspections.
We define the challenge of detecting anomalous regions between a reference non-anomalous set of images and a recapture of the scene with unaligned views as Scene Anomaly Detection (Scene AD). The recaptured scene will be referred to as the query scene and the corresponding set of images as the query set. Importantly, the query scene may or may not contain any anomalies. Additionally, what may constitute as an anomalous region is not available in most scenarios, thus providing supervision for supervised methods is usually unfeasible. Therefore, Scene AD requires methods that do not rely on any supervision for what may constitute as an anomalous region.
Prior unsupervised anomaly detection methods have not been designed to handle Scene AD and are typically developed for controlled industrial settings with constant camera positions. As a result, we find that state-of-the art anomaly detection methods perform poorly on the Scene AD task. Apart from the architectures themselves, we find that there are two main challenges that need to be overcome. First, there is no benchmark multi-object dataset available for Scene AD. As described earlier, existing AD benchmarks typically focus on a single view. Zhou et al. proposed a multi-pose AD [8] dataset, however this problem setting focuses on multiple poses of a single object and not that of an entire scene. Second, any solution to Scene AD must thus incorporate a mechanism for view-invariance.
Due to the limited generalizability of prior unsupervised anomaly detection (AD) methods in Scene AD, we propose OmniAD, a refinement of Reverse Distillation (RD) [9]. To detect anomalous region locations using RD, deviations in a student decoder must occur with respect to a teacher encoder. We postulate that expanding the effective receptive field (ERF) [10] of the student decoder will increase the deviations produced by the student decoder, which in turn will increase the sensitivity to anomalous regions. To expand the ERF, we incorporate student attention modules into the student decoder. We show quantitatively that the expansion of the ERF will, in turn, increase the performance of RD.
However, to achieve improved view-invariance, we introduce two new data augmentation methods. Building on the intuition that it is much easier to compare perfectly aligned pictures, rather than comparing misaligned ones, our data augmentation methods utilize novel view synthesis (NVS) techniques. Specifically, we employ Neural Radiance Fields (NeRF) [11], combined with a camera localization strategy, hierarchical localization (hloc) [12, 13], to generate novel views of the query scene that closely align with the non-anomalous views. We hypothesize that providing these generated aligned novel views will allow AD methods to improve generalizability. Although supervised methods for anomaly detection cannot leverage this data augmentation due to the lack of ground truth for anomalies, unsupervised methods can effectively use this approach for improved performance.
To properly evaluate and benchmark the effectiveness of these methods, a suitable dataset is required. Recognizing the lack of a comprehensive benchmark for Scene AD in real scenes with multiple objects, we introduce the ToyCity dataset. ToyCity is tailored to provide a robust real benchmark featuring multiple camera viewpoints and a range of anomalies across different setups. ToyCity’s unique structure presents a new and significant challenge to state-of-the-art AD models. We evaluate several AD methods, including OmniAD, using ToyCity as well as an object focused dataset, the Multi-pose Anomaly Detection dataset (MAD) [8].
To summarize our main contributions are:
  • Introduction of the novel computer vision task of Scene AD involving detecting anomalous regions between a reference non-anomalous set of images and a recapture of the scene with unaligned views. Results are presented to demonstrate the inability of existing anomaly detection methods to generalize effectively to the Scene AD task.
  • The ToyCity dataset specifically designed for Scene AD, addressing the need for suitable benchmarks for multi-object and multi-view problems in this domain.
  • Development of a new anomaly detection architecture termed OmniAD that better generalizes to Scene AD. Specifically, we introduce a novel Student Attention Module to expand the effective receptive field (ERF) of the reverse distillation (RD) architecture and demonstrate state-of-the art performance on the ToyCity and MAD datasets [8].
  • Two novel data augmentation techniques leveraging Neural Radiance Fields and hierarchical localization, resulting in a 40% increase in performance for pixel-wise localization of anomalous regions in OmniAD.
This paper is organized as follows: II provides a review of the relevant literature in the area of anomaly detection, neural radiance fields, and hierarchical localization. III describes the proposed methodology for data augmentation as well as OmniAD. IV discusses the structure of the ToyCity dataset, developed to evaluate AD methods for scene AD. V describes the experimental setup, evaluation metrics, and both a quantitative and qualitative analysis of the experimental results. Finally, VI is a conclusion to the paper, summarizing the main findings, followed by references.
II Related Work
Anomaly Detection. AD algorithms typically harness a collection of non-anomalous images , to establish a normative distribution. AD is a critical function in industrial visual inspection for identifying defects [14] and is also employed in healthcare for the detection of irregularities in medical images [15]. Prominent AD methods include CFLOW-AD [16], FastFlow [17], CFA [18], DRÆM [19], and Reverse Distillation (RD) [9].
Both CFLOW-AD and FastFlow utilize a flow framework [20] that models the distribution of normal image patches to detect anomalous patterns. Typically, these methods begin by extracting features from a pretrained CNN encoder for an input image patch. The feature vectors are then mapped to a multivariate Gaussian distribution through the flow framework. Anomalous images can then be detected because they tend to fit poorly to this learned distribution, highlighting their abnormality.
The CFA method uses a combination of a frozen pretrained CNN, a patch descriptor network that learns on-the-fly, and a memory bank for storing compressed features. CFA fine-tunes the pretrained CNN features through metric learning, which is based on a coupled hypersphere approach, clustering normal features tightly within a hypersphere’s boundary.
DRÆM operates with two main networks: a reconstructive network and a discriminative network. DRÆM generates synthetic anomalies that the reconstructive network then transforms back to normal. The discriminative network uses both the reconstructed normal image and the synthetic anomalous image to predict segmentation for the regions where the synthetic anomaly was introduced.
RD adopts a knowledge distillation strategy [21, 22], where a student decoder network learns the feature space representation from a teacher encoder network at various feature map scales. Unlike traditional approaches that feed the input image into both the student and the teacher networks, RD takes an autoencoder-like approach where the teacher encodes the input image, and the student decodes a fused multi-scale feature map output by the teacher. This feature map is a compressed representation of the teacher’s features across multiple scales.
Predominant datasets for benchmarking these methods, like MVTecAD [14] and VisA [23], typically employ a fixed overhead camera perspective, resulting in a uniform view of anomalies. Our research demonstrates that significantly different views can lead to substantial degradation in the pixel-level localization of anomalies. To combat these limitations, our method leverages a novel view synthesis method and camera localization techniques to generate corresponding non-anomalous images, thus facilitating improved anomaly detection. While there are existing methods for 3D anomaly detection, they generally require 3D data that is aligned with a reference model.
Neural Radiance Field NeRFs, as introduced by Mildenhall et al. [11], represent a significant advancement in the field of view synthesis. These fields use a collection of images with known camera poses to synthesize new viewpoints of a scene. Unlike conventional 3D Structure-from-Motion (SfM) methods that construct explicit geometric models, NeRFs model a scene’s volume through a neural network that predicts color and density for space points along camera rays. This technique enables the creation of photorealistic images from novel perspectives. Subsequent innovations like Nerfacto [24] and Instant-ngp [25] have further refined NeRFs, offering rapid training and real-time rendering capabilities. In the context of anomaly detection, we leverage NeRFs to synthesize non-anomalous images from diverse viewpoints, significantly improving the detection process.
Hierarchical Localization The process of hierarchical localization (hloc) [26, 13] has been developed to accurately determine a camera’s position and orientation in space, a task known as 6-Degree-of-Freedom (DoF) localization [26, 13, 27]. This process begins with the extraction of local features from a query image, utilizing methods such as SuperPoint [12], Sift [28], or DISK [29]. A 3D SfM model serves as a reference for these local features, against which the query image is matched using global descriptors like VLAD [30]. The Perspective-n-Point algorithm, in conjunction with RANSAC, then utilizes these correspondences to estimate the camera’s pose. This pose is critical in enabling NeRFs to generate non-anomalous images that correspond to those containing anomalies, thus in theory providing a more robust dataset for anomaly detection algorithms. However as we shall show an improved AD method is needed to fully utilize these anomaly aligned images.
III Method
An overview of our AD methodology for Scene AD can be seen in Figure2. Scene AD consists of non-anomalous views and query views. We generate an SfM model using non-anomalous views to acquire camera poses. Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation as shown in Figure3. For the first strategy, we interpolate between the non-anomalous camera poses. The second strategy for generating useful camera views for training is to utilize hloc to acquire the camera poses of query views. We then use NVS with the new camera poses to augment our non-anomalous views for training our proposed method OmniAD. The following subsections describes each of these steps in more detail.
Figure 2: Overview of our methodology is depicted here. We utilize the non-anomalous views from a Scene AD to generate a non-anomalous SfM model. A novel view selection strategy utilizes anomaly views and the SfM model to generate novel views to augment the non-anomalous dataset. OmniAD is then trained on the augmented non-anomalous dataset for anomaly detection.
Figure 3: Overview of our novel view selection strategy. Utilizing the non-anomalous dataset, we generate a structure from motion (SfM) 3D model and accompanying camera pose for each image. Leveraging these non-anomalous camera poses, we design a trajectory that connects them, allowing for interpolation. Subsequently, we apply a NVS technique to generate images that present unique perspectives along the interpolated path. The anomaly dataset can also be used to determine more novel views.
III-A Novel View Selection Strategy
To generate new non-anomalous camera poses for NVS, we propose and evaluate two different approaches, namely Interpolated Non-anomalous Views (INV) and Query-aligned non-anomalous views (QANV).
III-A1 Interpolated Non-anomalous Views
Figure 4: The provided sequence illustrates camera interpolation, where the pose transitions from the starting image to the ending image, are used to synthesize intermediate views. These interpolated views, not present in the original dataset, bridge the gap between the start and end poses, introducing new perspectives to the dataset.
Given a set of non-anomalous images an associated camera poses $T$ can be calculated using SfM. The calculated $T$ can be interconnected to form a trajectory. A simple approach for this interconnection is a greedy algorithm that selects a starting camera position and sequentially moves to the nearest neighboring camera position. Specifically, for our experiments, we interpolate 12 intermediate camera poses between consecutive original poses in $T$, thereby enhancing the dataset with interpolated views that increase the robustness of our AD model to various camera angles. An example of the camera interpolation for ToyCity can be seen in Figure4.
III-A2 Query-Aligned Non-anomalous Views
Here we propose a method to generate images of the baseline scene, but using poses from the anomaly scene. Our method employs hloc to accurately estimate the camera pose for a query anomaly image. This necessitates an accurate reconstruction of the baseline scene that is provided by the baseline SfM model. Hloc provides a variety of options for local feature detection, matching, and image similarity measurement. We opted for a combination of SuperPoint trained on Aachen and SuperGlue [12, 31], as it demonstrated effective performance on our datasets on average. Nonetheless, the flexibility to switch to other methods is inherent in our approach, enabling adaptation to specific dataset characteristics.
III-B Novel View Synthesis
We selected Nerfacto [24] due to its balanced performance in training speed and quality compared to other methods, along with its user-friendly interface, which includes an interactive viewer. Although our method is agnostic to the specific NVS technique used, the aforementioned benefits made Nerfacto a compelling choice for this work. Continual advancements in this field, like the advent of 3D Gaussian Splattering[32], present opportunities for further enhancing rendering time and quality.
Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure3. Example camera interpolated images can be seen in Figure3, and example images using QANV can be seen in Figure3
III-C OmniAD
We develop OmniAD to leverage our proposed augmentation techniques and surpass the performance of state-of-the-art AD methods when applied to Scene AD. OmniAD significantly enhances the RD framework through two significant updates, (i) integrating a ResNeXt backbone in the teacher encoders[33] (ii) and adding novel student attention modules, $A^{i}$. We found a combination of the ResNext with the student attention module to offer a performance increase of up to 40%, while the training and inference time cost for OmniAD is similar to RD.
Figure 5: OmniAD’s architecture mirrors RD’s, featuring a frozen teacher encoder that channels features through a bottleneck module into a student decoder. We enhance this setup by integrating student attention modules $A^{1}$, $A^{2}$, and $A^{3}$, bolstering the architecture’s proficiency in utilizing Query-aligned non-anomalous views and Interpolated Baseline Views for improved effectiveness.
Student attention modules $A^{i}$ are introduced after each of the Student Decoder networks $D^{i}$ of the RD architecture. These modules are designed to bolster OmniAD’s sensitivity to anomalous samples. An illustrative overview of the OmniAD architecture is presented in Figure5. $A^{i}$ is based on the U-Net architecture and consists of alternating layers of 3x3 convolutions with self attention layers and a bottleneck layer consisting of several self attention layers as shown in Figure6. This architecture was designed to expand the Effective Receptive Field (ERF) of the Student Decoder which in turn would increase the variations in the output features in response to an anomaly. We validate this hypothesis and visualize the larger ERF in Figure7. This means there is a larger perturbation of features when an anomalous region is introduced into the model, making it more sensitive and responsive to such anomalies. This increased sensitivity is attributed to the improved performance of OmniAD. We conducted experiments to validate this hypothesis and demonstrate the efficacy of $A^{i}$.
Figure 6: Student attention module.
(a) with $A^{I}$
(b) w/o $A^{I}$
Figure 7: ERF of OmniAD with and without the student attention module. (a) with $A^{I}$. (b) w/o $A^{I}$.
IV ToyCity Dataset
The ToyCity dataset addresses the need for a benchmark dataset consisting of real images of multiple objects for Scene AD. Unlike MAD, ToyCity consists of (i) substantially more real images and (ii) is not centered around a single object. ToyCity features scenes of a toy city with buildings, signs, roads, trees, and miscellaneous items. The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure8. Each setup also contains different anomalies, as shown in Figure8. Additoinally, each anomaly image has at most one anomaly. The dataset includes the manually annotated ground truth for anomaly segmentation. Table I provides an overview of the number of images released for ToyCity. The dataset’s v1 is the base variant of all four scenes without any augmentations, v2 includes Quer-Aligned Novel Views (QANV) images rendered using Nerfacto, v3 incorporates augmentations using Interpolated Non-Anomalous Views (INV), and v4 combines the data of both v2 and v3, these variations are presented in Table II. ToyCity thus offers a new benchmarking challenge to current state-of-the-art anomaly detection models.
Figure 8: ToyCity consists of four configurations as shown here. We present an example baseline image for each as well as an example of one of several anomalies used. A zoomed comparison is also presented that shows the anomaly highlighted with a red outline.
TABLE I: Number of non-anomalous and anomalous images in each of the ToyCity dataset scenes, and the number of NVS images generated from each of the proposed methods QANV and IBV
Name Non-Anomalous Anomalous QANV IBV
ToyCity 1 316 319 319 3781
ToyCity 2 341 308 308 3773
ToyCity 3 307 317 317 3673
ToyCity 4 314 314 314 3757
V Experiments
We evaluate several AD methods on ToyCity and MAD to test the localization ability of each method. We could only utilize a subset of the MAD dataset that consisted of real images compatible with COLMAP for Structure from Motion (SfM). Each AD method is evaluated quantitatively using the pixel-level $F_{1}$ score and AUROC. An ablation study follows to show the significance of the student attention modules on RD. Finally a qualitative comparison between OmniAD and RD for anomaly segmentation is presented.
V-A Metrics
To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$ Score as described by Equation 1. The $F_{1}$ score excels in handling class imbalances by merging precision and recall into one comprehensive metric. This balance is critical for datasets such as ToyCity and MAD, where anomaly regions typically constitute a small fraction of the total image area. Examples of the relatively small anomaly regions within ToyCity can be seen in Figure8, underscoring the necessity for such a balanced evaluation metric. The imbalance of anomaly regions renders other metrics like accuracy or the area under the receiver operating characteristic curve (AUROC) as misleading.
$F_{1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}$ (1)
The reason accuracy is deceptive is that a high accuracy metric can be achieved with a simplistic approach that labels all pixels as non-anomalous. This approach would yield numerous true negatives and few false positives. A similar effect can be observed with the AUROC, which may show a more optimistic estimate of model performance. However, it does not accurately reflect the model’s ability to precisely localize anomalies, primarily due to the abundance of true negatives. Thus we provide both the $F_{1}$ score and AUROC.
V-B Implementation Details
To allow for fair comparison between all anomaly detection methods we use the anomalib [34] library implementation of each anomaly detection method. A fixed random seed of 42 was used for all methods. All methods were allowed to train for a maximum of 100 epochs and the max achieved $F_{1}$ score was selected as the best model. All methods converge well before 100 epochs were completed ensuring fair training. Additionally methods were trained with the same batch size. Other hyper-parameters were left as default values provided by anomalib for each anomaly detection method.
V-C Ablation Studies
V-C1 OmniAD architecture
OmniAD has two major changes from RD. First we change the backbone to ResNeXt and second we introduced branching student attention modules $A^{i}$. To test if our changes are significant we test OmniAD with and without ResNeXt (OmniAD-R) and without $A^{i}$ (OmniAD-$A^{i}$) against RD. As shown in the $F_{1}$ scores on ToyCity Tables III and MAD Tables V we can see a consistent performance increase when $A^{i}$ is used.
V-C2 NVS Augmentation
We test the efficacy of our NVS augmentation strategies by evaluating our metrics with and without each of the augmentations. Our hypothesis that generating novel views of the our non-anomalous scene similar to the view point of the query scene will make it easier for OmniAD to identify anomalies is verified through four different variations listed in Table II.
V-D Quantitative Evaluation
In our assessment of diverse AD methods on the ToyCity datasets, we observe a consistent pattern of performance increase OmniAD, OmniAD without ResNext (OmniAD w/o R), and RD when a novel view selection strategy is employed for $F_{1}$ scores on ToyCity and MAD. A similar trend can also be seen for AUROC on both ToyCity and MAD Tables IV,VI. Additional experiments on different scenes is provided in the supplemental material.
The RD with ResNext model and the canonical RD methodology exhibit similar enhancement trends, yet they exhibit smaller gains in comparison to our model. The shift from v1 to v2, where anomaly aligned images are introduced, is particularly telling of the significant enhancements in the models’ ability to localize anomalies. Other AD methods such as CFLOW, FastFlow, CFA, and DRÆM show a range of responses to the augmentations, with $F_{1}$ Scores experiencing minimal improvements, remaining unchanged, or occasionally declining with the introduction of more complex augmentations. This suggests a potential lack of effective utilization of augmented data by these methods in contrast to OmniAD and RD.
As mentioned metrics such as pixel-level AUROC can often show very high numbers for datasets such as ToyCity and MAD due to the low number of pixels that may be an anomaly. This can be seen for both ToyCity Table IV and MAD Table VI. Most AD methods will achieve an AUROC very close to 1 while still offering low $F_{1}$ scores.
TABLE II: Four variations are tested for each of the scenes present in ToyCity. Variation v1 is un-augmented, v2 is with the Query-Aligned Views, v3 is with the interpolated Non-Anomalous views, and v4 is the combination of Query-Aligned Baseline views and Interpolated Non-Anomalous views.
Variation Non-Anomalous Anomalous QANV INV
v1 - -
v2 -
v3 -
v4
V-E Qualitative Evaluation
Figure9 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, RD, which utilizes a ResNeXt backbone. A threshold was selected for each method that maximizes their pixel-level $F_{1}$ score. Three primary failure scenarios are evident. First, as depicted in the ToyCity 1 and 4 anomaly instances, RD may erroneously identify arbitrary areas as anomalous. Secondly, RD has a propensity to overlook entire anomalous regions, a shortcoming exemplified by the ToyCity 2 anomaly. Lastly, RD often overestimates the extent of an anomaly, erroneously marking larger areas than necessary. Similar defects in RD can also be seen from the MAD dataset in which it misses an anomaly or overestimates an anomalous region. Additional samples from OmniAD for each dataset is presented in Figure 10, showing a heat map of the anomaly score as well as the thresholded segmentation.
Figure 9: Segmentation performance comparison between OmniAD and RD with ResNeXt backbone across four ToyCity anomalies and 2 objects from MAD.We illustrating three primary failure patterns: misidentifying of non-anomalous areas (ToyCity 1,4m and MAD Pig), overlooking anomalies (ToyCity 2 and Mad Puppy), and overestimation of anomaly areas (ToyCity 3 and MAD Pig).
Figure 10: Additional segmentations and anomaly scorces from OmniAD.
TABLE III: Comparison of Pixel $F_{1}$ scores on ToyCity 1, 2, 3, and 4
Method ToyCity 1 ToyCity 2 ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.432 0.450 0.497 0.496 0.348 0.599 0.491 0.577 0.281 0.384 0.319 0.366 0.379 0.541 0.390 0.523
OmniAD-R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
OmniAD-$A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
RD 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
CFLOW 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
FastFlow 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
CFA 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
DRÆM 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
TABLE IV: Comparison of Pixel AUROC scores on ToyCity 1, 2, 3, and 4
Method ToyCity 1 ToyCity 2 ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.990 0.996 0.999 0.999 0.991 0.998 0.999 0.999 0.979 0.993 0.997 0.999 0.982 0.993 0.995 0.998
OmniAD-R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
OmniAD-$A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
RD 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
CFLOW 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
FastFlow 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
CFA 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
DRÆM 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
TABLE V: Comparison of Pixel $F_{1}$ scores on MAD-Real [8] of Puppy, Pig, and Elephant
Method Puppy Pig Elephant
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.672 0.714 0.680 0.635 0.334 0.395 0.471 0.524 0.341 0.368 0.374 0.328
OmniAD-R 0.613 0.677 0.621 0.623 0.302 0.346 0.454 0.441 0.312 0.204 0.297 0.284
OmniAD-$A^{i}$ 0.634 0.646 0.377 0.363 0.306 0.286 0.343 0.254 0.275 0.298 0.248 0.173
RD 0.616 0.628 0.520 0.199 0.219 0.203 0.302 0.260 0.262 0.236 0.170 0.192
CFLOW 0.496 0.466 0.484 0.579 0.162 0.142 0.150 0.147 0.309 0.163 0.214 0.209
FastFlow 0.646 0.651 0.575 0.598 0.066 0.087 0.043 0.058 0.297 0.285 0.282 0.297
CFA 0.568 0.576 0.594 0.067 0.294 0.232 0.048 0.029 0.189 0.151 0.117 0.139
DRÆM 0.042 0.016 0.022 0.018 0.040 0.041 0.072 0.110 0.020 0.041 0.037 0.028
TABLE VI: Comparison of Pixel AUROC scores on MAD-Real [8] of Puppy, Pig, and Elephant
Method Puppy Pig Elephant
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.999 0.999 0.999 0.999 0.988 0.993 0.999 0.999 0.941 0.968 0.973 0.971
OmniAD-R 0.999 0.993 0.999 0.999 0.988 0.983 0.999 0.999 0.924 0.892 0.933 0.928
OmniAD-$A^{i}$ 0.999 0.999 0.998 0.999 0.983 0.988 0.997 0.997 0.948 0.946 0.966 0.963
RD 0.999 0.999 0.999 0.999 0.981 0.982 0.998 0.998 0.936 0.929 0.951 0.956
CFLOW 0.996 0.996 0.999 0.999 0.962 0.967 0.993 0.990 0.945 0.906 0.947 0.953
FastFlow 0.999 0.998 0.999 0.999 0.868 0.896 0.965 0.966 0.862 0.914 0.934 0.932
CFA 0.998 0.998 0.999 0.994 0.977 0.976 0.986 0.979 0.916 0.927 0.934 0.942
DRÆM 0.819 0.359 0.514 0.524 0.720 0.571 0.684 0.796 0.601 0.496 0.559 0.519
VI Conclusion
Our study introduced and addressed a specific anomaly detection (AD) problem we term Scene AD. Scene AD deals with pixel-level AD between a non-anomalous scene and an anomalous scene, where a misalignment between non-anomalous and anomalous views is expected, as is often the case in inspections of uncontrolled environments. To challenge and measure the effectiveness of state-of-the-art anomaly detection for Scene AD, we introduced the ToyCity dataset. This new benchmark focuses on viewpoint variance in scenes of multiple objects and serves as a testbed for our new method OmniAD. We demonstrated through extensive comparisons that OmniAD outperforms other AD methods when applied to the ToyCity dataset, establishing a new state-of-the-art for this task. Particularly, we found that the incorporation of a student attention module increases the efficacy of reverse distillation for pixel-level AD by expanding the effective receptive filed (ERF). Our experiments also demonstrated the benefits of novel view synthesis and camera localization for dataset augmentation in enhancing the robustness and accuracy of OmniAD. We hope that our work will catalyze further advancements, spurring the creation of advanced anomaly detection techniques capable of managing the complexities associated with varying viewpoints in 3D scenes.

### v2

@@
-View-Invariant Pixelwise Anomaly Detection in Multi-object Scenes with Adaptive View Synthesis
+OmniAD: View-Invariant Unsupervised Scene Pixelwise Anomaly Detection with Adaptive View Synthesis
 Abstract
-The inspection and monitoring of infrastructure assets typically requires identifying visual anomalies in scenes periodically photographed over time.
-Images collected manually or with robots such as unmanned aerial vehicles from the same scene at different instances in time are typically not perfectly aligned.
-Supervised segmentation methods can be applied to identify known problems, but unsupervised anomaly detection approaches are required when unknown anomalies occur.
-Current unsupervised pixel-level anomaly detection methods have mainly been developed for industrial settings where the camera position is known and constant.
-However, we find that these methods fail to generalize to the case when images are not perfectly aligned.
-We term the problem of unsupervised anomaly detection between two such imperfectly aligned sets of images as Scene Anomaly Detection (Scene AD).
-We present a novel network termed OmniAD to address the Scene AD problem posed.
-Specifically, we refine the anomaly detection method reverse distillation to achieve a 40% increase in pixel-level anomaly detection performance.
-The network’s performance is further demonstrated to improve with two new data augmentation strategies proposed that leverage novel view synthesis and camera localization to improve generalization.
-We validate our approach with qualitative and quantitative results on a new dataset, ToyCity, the first Scene AD dataset with multiple objects, as well as on the established single object-centric dataset, MAD.
-https://drags99.github.io/OmniAD/
-Index Terms:
-Anomaly Detection, Unsupervised Learning, Camera Localization
-I Introduction
-Figure 1: In Scene AD, we observe variations in camera positions between non-anomalous views (represented in green), and query views which may have anomalies in them (represented in red).
-In contrast, traditional AD methods typically feature aligned views between the non-anomalous views and anomalous views.
-The inspection and monitoring of infrastructure assets typically requires identifying visual anomalies in scenes periodically photographed over time [1, 2, 3, 4, 5].
-Images collected manually or with robots such as unmanned aerial vehicles from the same scene at different instances in time are typically not perfectly aligned.
-For instance, the United States Navy places significant emphasis on the monitoring and detection of anomalies during inspections of their ships, which is critical for maintaining security and operational readiness and safety [6, 7].
-Similarly, bridge inspectors can leverage this technology to identify parts of the bridge that are changing or behaving abnormally.
-This information is vital for predicting future performance and for maintenance decision making.
-In each of these examples, new image sets are acquired at every inspection but typically unaligned with previous inspections.
-We define the challenge of detecting anomalous regions between a reference non-anomalous set of images and a recapture of the scene with unaligned views as Scene Anomaly Detection (Scene AD).
-The recaptured scene will be referred to as the query scene and the corresponding set of images as the query set.
-Importantly, the query scene may or may not contain any anomalies.
-Additionally, what may constitute as an anomalous region is not available in most scenarios, thus providing supervision for supervised methods is usually unfeasible.
-Therefore, Scene AD requires methods that do not rely on any supervision for what may constitute as an anomalous region.
-Prior unsupervised anomaly detection methods have not been designed to handle Scene AD and are typically developed for controlled industrial settings with constant camera positions.
-As a result, we find that state-of-the art anomaly detection methods perform poorly on the Scene AD task.
-Apart from the architectures themselves, we find that there are two main challenges that need to be overcome.
-First, there is no benchmark multi-object dataset available for Scene AD.
-As described earlier, existing AD benchmarks typically focus on a single view.
-Zhou et al.
-proposed a multi-pose AD [8] dataset, however this problem setting focuses on multiple poses of a single object and not that of an entire scene.
-Second, any solution to Scene AD must thus incorporate a mechanism for view-invariance.
-Due to the limited generalizability of prior unsupervised anomaly detection (AD) methods in Scene AD, we propose OmniAD, a refinement of Reverse Distillation (RD) [9].
-To detect anomalous region locations using RD, deviations in a student decoder must occur with respect to a teacher encoder.
-We postulate that expanding the effective receptive field (ERF) [10] of the student decoder will increase the deviations produced by the student decoder, which in turn will increase the sensitivity to anomalous regions.
-To expand the ERF, we incorporate student attention modules into the student decoder.
-We show quantitatively that the expansion of the ERF will, in turn, increase the performance of RD.
-However, to achieve improved view-invariance, we introduce two new data augmentation methods.
-Building on the intuition that it is much easier to compare perfectly aligned pictures, rather than comparing misaligned ones, our data augmentation methods utilize novel view synthesis (NVS) techniques.
-Specifically, we employ Neural Radiance Fields (NeRF) [11], combined with a camera localization strategy, hierarchical localization (hloc) [12, 13], to generate novel views of the query scene that closely align with the non-anomalous views.
-We hypothesize that providing these generated aligned novel views will allow AD methods to improve generalizability.
-Although supervised methods for anomaly detection cannot leverage this data augmentation due to the lack of ground truth for anomalies, unsupervised methods can effectively use this approach for improved performance.
-To properly evaluate and benchmark the effectiveness of these methods, a suitable dataset is required.
-Recognizing the lack of a comprehensive benchmark for Scene AD in real scenes with multiple objects, we introduce the ToyCity dataset.
-ToyCity is tailored to provide a robust real benchmark featuring multiple camera viewpoints and a range of anomalies across different setups.
-ToyCity’s unique structure presents a new and significant challenge to state-of-the-art AD models.
-We evaluate several AD methods, including OmniAD, using ToyCity as well as an object focused dataset, the Multi-pose Anomaly Detection dataset (MAD) [8].
-To summarize our main contributions are:
-  • Introduction of the novel computer vision task of Scene AD involving detecting anomalous regions between a reference non-anomalous set of images and a recapture of the scene with unaligned views.
-Results are presented to demonstrate the inability of existing anomaly detection methods to generalize effectively to the Scene AD task.
-  • The ToyCity dataset specifically designed for Scene AD, addressing the need for suitable benchmarks for multi-object and multi-view problems in this domain.
-  • Development of a new anomaly detection architecture termed OmniAD that better generalizes to Scene AD.
-Specifically, we introduce a novel Student Attention Module to expand the effective receptive field (ERF) of the reverse distillation (RD) architecture and demonstrate state-of-the art performance on the ToyCity and MAD datasets [8].
-  • Two novel data augmentation techniques leveraging Neural Radiance Fields and hierarchical localization, resulting in a 40% increase in performance for pixel-wise localization of anomalous regions in OmniAD.
-This paper is organized as follows: II provides a review of the relevant literature in the area of anomaly detection, neural radiance fields, and hierarchical localization.
-III describes the proposed methodology for data augmentation as well as OmniAD.
-IV discusses the structure of the ToyCity dataset, developed to evaluate AD methods for scene AD.
-V describes the experimental setup, evaluation metrics, and both a quantitative and qualitative analysis of the experimental results.
-Finally, VI is a conclusion to the paper, summarizing the main findings, followed by references.
-II Related Work
+Visual anomaly detection in the built environment can be a valuable tool for applications such as infrastructure assessment, construction monitoring, security surveillance, and urban planning.
+Anomaly detection approaches are typically unsupervised and work by detecting deviations from an expected state where no assumptions are made exact type of deviation.
+Unsupervised pixel-level anomaly detection methods have been developed to successfully recognize and segment anomalies; however, existing techniques are designed for industrial settings with a fixed camera position.
+In the built environment, images may be periodically collected by a moving camera manually or by aerial or ground vehicles and utilized for decision making.
+The camera pose between successive collections may vary widely voiding a fundamental assumption in existing anomaly detection approaches.
+To address this gap, we introduce the problem of Scene Anomaly Detection (Scene AD), where the goal is to detect anomalies from two sets of images: one set without anomalies and one set that may or may not contain anomalies.
+No labeled semantic segmentation data are provided for training.
+We propose a novel network, OmniAD, to tackle Scene AD by refining the reverse distillation anomaly detection method, leading to a 40% improvement in pixel-level anomaly detection.
+Additionally, we introduce two new data augmentation strategies that leverage novel view synthesis and camera localization to enhance generalization.
+We evaluate our approach both qualitatively and quantitatively on a new dataset, ToyCity the first Scene AD dataset featuring multiple objects as well as on the established single object centric dataset, MAD.
+Our method demonstrates marked improvement over baseline approaches, paving the way for robust anomaly detection in scenes with real-world camera pose variations commonly observed in the built environment.
+Project Page: https://drags99.github.io/OmniAD/
+keywords:
+Anomaly Detection , Unsupervised Learning , Camera Localization , Neural Radiance Fields
+[…]
+[1] organization=Department of Electrical and Computer Engineering, University of Houston, addressline=4226 Martin Luther King Blvd., city=Houston, state=Texas, postcode=77204, country=USA
+[…]
+[2] organization=Department of Civil and Environmental Engineering, University of Houston, addressline=4226 Martin Luther King Blvd., city=Houston, state=Texas, postcode=77204, country=USA
+1 Introduction
+Figure 1: In Scene AD, we observe variations in camera positions between non-anomalous views (green), and query views which may contain anomalies (red).
+In contrast, traditional AD methods typically feature aligned views between the non-anomalous and anomalous images.
+For many applications in the built environment —ranging from infrastructure assessment and construction monitoring to security surveillance and urban planning—detecting visual anomalies in scenes captured over time is a critical task  [1, 2, 3, 4, 5].
+These anomalies may arise from structural changes, newly introduced objects, unexpected modifications in appearance, or more generally, anything that deviates from an established non-anomalous reference.
+However, images acquired at different instances, whether manually or autonomously, typically suffer from misalignment due to variations in camera pose.
+For example, bridge inspectors may assess changes to a structure following a disaster, self-driving car systems might detect modifications to traffic signs, and the U.S.
+Military places strong emphasis on monitoring infrastructure for anomalies critical to operational readiness [6, 7, 8, 9, 10, 11, 12].
+In all such cases, each new set of images is unaligned with previous captures, complicating anomaly identification.
+We define the task of Scene Anomaly Detection (Scene AD) as detecting anomalous regions in a query set of images, potentially containing no anomalies at all, against a non-anomalous reference set.
+Labeled examples of what qualifies as an anomaly are unavailable, rendering supervised segmentation methods impractical.
+Additionally anomalies are often too infrequent or unpredictable to compile sufficient labeled training data for every potential anomaly type, rendering fully supervised models infeasible.
+As a result, unsupervised anomaly detection (AD) approaches are a feasible alternative.
+However, these AD methods were originally crafted for controlled industrial scenarios which commonly assume fixed camera views and struggle with the viewpoint variations of real-world Scene AD.
+Two main challenges emerge.
+First, there is no suitable multi-object, multi-view benchmark for Scene AD: most existing AD datasets capture a single view or focus on a single object [13].
+Second, any robust Scene AD solution must tolerate significant camera misalignment and still detect newly introduced or changed regions.
+To address these gaps, we propose OmniAD, a refinement of the Reverse Distillation (RD) framework [14].
+RD detects anomalies by comparing a student decoder’s reconstruction with a teacher encoder’s embedding.
+We hypothesize that broadening the student decoder’s effective receptive field (ERF) [15] will improve sensitivity to anomalous regions.
+Accordingly, we incorporate student attention modules to expand the ERF and enhance pixel-level anomaly detection.
+Moreover, to manage viewpoint differences in real scenes, we introduce two data augmentation strategies leveraging Neural Radiance Fields (NeRF) [16] and hierarchical localization (hloc) [17, 18].
+These techniques generate novel, more closely aligned views of the query scene, aiding unsupervised methods—such as OmniAD—in detecting anomalies without labeled examples.
+Since no comprehensive benchmark exists for this multi-object, multi-view setting, we introduce the ToyCity dataset.
+ToyCity consists of scenes with multiple objects, varied camera viewpoints, and different anomalies.
+We evaluate OmniAD and several baseline methods on ToyCity as well as on the single-object-centric Multi-pose Anomaly Detection dataset (MAD) [13], demonstrating notable performance gains.
+Our main contributions are:
+  • 1.
+    Scene AD Task: Formal introduction of Scene Anomaly Detection, highlighting the challenge of detecting anomalies in unaligned multi-view imagery.
+  • 2.
+    ToyCity Dataset: A new multi-object, multi-view dataset designed for benchmarking Scene AD methods in realistic settings.
+  • 3.
+    OmniAD Architecture: A refined Reverse Distillation approach with a proposed Student Attention Module to expand the ERF, achieving state-of-the-art performance on ToyCity and MAD [13] and validated with ablation studies.
+  • 4.
+    NeRF-based Data Augmentation Methods: Two novel techniques using NeRF and hierarchical localization for improved view-invariance, yielding a 40% increase in anomaly localization accuracy.
+The rest of the paper is organized as follows: Section 2 reviews related work on anomaly detection, neural radiance fields, and camera localization.
+Section 3 introduces the proposed OmniAD architecture and novel data augmentation strategies.
+Section 4 presents the ToyCity dataset.
+Section 5 details the experimental setup, metrics, and results.
+Section 6 discusses the limitations of our dataset and method.
+Finally, Section 7 concludes the paper.
+2 Related Work
 Anomaly Detection.
@@
 However as we shall show an improved AD method is needed to fully utilize these anomaly aligned images.
-III Method
-An overview of our AD methodology for Scene AD can be seen in Figure2.
+3 Method
+An overview of our AD methodology for Scene AD can be seen in Figure  2.
 Scene AD consists of non-anomalous views and query views.
 We generate an SfM model using non-anomalous views to acquire camera poses.
-Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation as shown in Figure3.
+Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation as shown in Figure  3.
 For the first strategy, we interpolate between the non-anomalous camera poses.
@@
 The anomaly dataset can also be used to determine more novel views.
-III-A Novel View Selection Strategy
+3.1 Novel View Selection Strategy
 To generate new non-anomalous camera poses for NVS, we propose and evaluate two different approaches, namely Interpolated Non-anomalous Views (INV) and Query-aligned non-anomalous views (QANV).
-III-A1 Interpolated Non-anomalous Views
+3.1.1 Interpolated Non-anomalous Views
 Figure 4: The provided sequence illustrates camera interpolation, where the pose transitions from the starting image to the ending image, are used to synthesize intermediate views.
@@
 Specifically, for our experiments, we interpolate 12 intermediate camera poses between consecutive original poses in $T$, thereby enhancing the dataset with interpolated views that increase the robustness of our AD model to various camera angles.
-An example of the camera interpolation for ToyCity can be seen in Figure4.
-III-A2 Query-Aligned Non-anomalous Views
+An example of the camera interpolation for ToyCity can be seen in Figure  4.
+3.1.2 Query-Aligned Non-anomalous Views
 Here we propose a method to generate images of the baseline scene, but using poses from the anomaly scene.
@@
 Nonetheless, the flexibility to switch to other methods is inherent in our approach, enabling adaptation to specific dataset characteristics.
-III-B Novel View Synthesis
+3.2 Novel View Synthesis
 We selected Nerfacto [29] due to its balanced performance in training speed and quality compared to other methods, along with its user-friendly interface, which includes an interactive viewer.
@@
 Continual advancements in this field, like the advent of 3D Gaussian Splattering[37], present opportunities for further enhancing rendering time and quality.
-Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure3.
-Example camera interpolated images can be seen in Figure3, and example images using QANV can be seen in Figure3
-III-C OmniAD
-We develop OmniAD to leverage our proposed augmentation techniques and surpass the performance of state-of-the-art AD methods when applied to Scene AD.
-OmniAD significantly enhances the RD framework through two significant updates, (i) integrating a ResNeXt backbone in the teacher encoders[33] (ii) and adding novel student attention modules, $A^{i}$.
-We found a combination of the ResNext with the student attention module to offer a performance increase of up to 40%, while the training and inference time cost for OmniAD is similar to RD.
+Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure  3.
+Example camera interpolated images can be seen in Figure  3, and example images using QANV can be seen in Figure  3
+3.3 OmniAD
+Our goal in Scene AD is to handle significant camera misalignment while detecting anomalies at a pixel level.
+Although Reverse Distillation (RD) [14] provides a robust teacher student mechanism for fixed view scenarios, our preliminary tests revealed its limitations in multi-view scenarios.
+Backbone Exploration.
+We first explored different backbones to replace RD’s default encoder, focusing on architectures that could accommodate the diverse viewpoints.
+Many alternatives struggled to converge or offered minimal gains, but ResNeXt [38] consistently yielded better multi-scale representations and improved performance compared to simpler encoders.
+In principle, any sufficiently expressive architecture including fully attention-based backbones could be substituted for the encoder.
+However, given the ever-growing variety of such networks, we restricted our exploration to widely adopted backbones and found that ResNeXt provided the most consistent gains for our multi-view setting.
+Receptive Field Constraints.
+Further analysis indicated that RD’s student decoder had a relatively small effective receptive field (ERF) [15], causing it to overlook anomalies.
+This shortcoming proved critical in multi-view settings, where anomalies could appear in various positions or at different scales.
+Rationale of ERF Expansion for Anomaly Sensitivity.
+Let $\mathcal{I}$ denote the image domain (e.g., $H\times W$ pixels), and let $x\in\mathbb{R}^{C\times H\times W}$ be an input image.
+Suppose a teacher network $T$ extracts multi-scale features
+$T(x)\;=\;\bigl{(}T^{1}(x),T^{2}(x),\dots,T^{L}(x)\bigr{)},$
+where $T^{l}(x)\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the $l$-th feature map.
+A student network $S$ decodes these teacher features to produce a reconstruction $\hat{x}=(S\circ T)(x)$.
+In Reverse Distillation [14], any mismatch between $x$ and $\hat{x}$ can indicate an anomaly, since $S$ is trained on non-anomalous (normal) data alone.
+Effective Receptive Field (ERF). For each output location $(i,j)$ in $\hat{x}$, define its effective receptive field as
+$\mathrm{ERF}_{ij}(S\circ T)\;=\;\bigl{\{}\,(u,v)\in\mathcal{I}\;\big{|}\;\tfrac{\partial\hat{x}_{ij}}{\partial x_{uv}}\neq 0\bigr{\}}.$
+In standard RD, convolutions and limited skip connections yield a relatively localized ERF.
+If a true anomaly lies outside $\mathrm{ERF}_{ij}$ for many $(i,j)$, it may not sufficiently alter $\hat{x}$, causing the anomaly to go undetected.
+Impact of ERF Expansion.
+Let $A\subset\mathcal{I}$ denote the anomaly region, i.e., pixels in $x$ that deviate from normal.
+By adding student attention modules (Fig.
+6), the new student $S^{\prime}$ broadens $\mathrm{ERF}_{ij}(S^{\prime}\circ T)$:
+$\bigl{|}\mathrm{ERF}_{ij}(S^{\prime}\circ T)\bigr{|}\;\;\geq\;\;\bigl{|}\mathrm{ERF}_{ij}(S\circ T)\bigr{|}.$
+Consequently, for any anomalous pixel $(u,v)\in A$, there are more $(i,j)$ in the reconstructed output $\hat{x}$ for which $(u,v)\in\mathrm{ERF}_{ij}(S^{\prime}\circ T)$.
+In other words, the anomaly influences a broader swath of the reconstruction, resulting in larger or more easily detectable errors.
+Reverse Distillation Anomaly Score.
+As in conventional RD, we define an anomaly score per pixel $(i,j)$ by measuring discrepancy between $\hat{x}_{ij}$ and $x_{ij}$.
+When the ERF is small, an anomaly’s influence might be confined or even “explained away” by the network averaging out the region, yielding low detection sensitivity.
+With an expanded ERF, the anomaly is more likely to cause a significant mismatch for at least one output location $(i,j)$, thereby increasing detection rates.
+Attention Modules for the Student Decoder.
+To address the ERF limitation, we introduced additional attention layers, denoted $A^{i}$, after each decoder block.
+These modules expand the ERF by allowing the network to aggregate information from spatially distant regions, thus capturing subtle or pose-dependent anomalies.
+Figure 6 shows the basic layout of each attention module, incorporating interleaved 3x3 convolutions and self-attention blocks for both local and global feature integration.
 Figure 5: OmniAD’s architecture mirrors RD’s, featuring a frozen teacher encoder that channels features through a bottleneck module into a student decoder.
 We enhance this setup by integrating student attention modules $A^{1}$, $A^{2}$, and $A^{3}$, bolstering the architecture’s proficiency in utilizing Query-aligned non-anomalous views and Interpolated Baseline Views for improved effectiveness.
-Student attention modules $A^{i}$ are introduced after each of the Student Decoder networks $D^{i}$ of the RD architecture.
-These modules are designed to bolster OmniAD’s sensitivity to anomalous samples.
-An illustrative overview of the OmniAD architecture is presented in Figure5.
-$A^{i}$ is based on the U-Net architecture and consists of alternating layers of 3x3 convolutions with self attention layers and a bottleneck layer consisting of several self attention layers as shown in Figure6.
-This architecture was designed to expand the Effective Receptive Field (ERF) of the Student Decoder which in turn would increase the variations in the output features in response to an anomaly.
-We validate this hypothesis and visualize the larger ERF in Figure7.
-This means there is a larger perturbation of features when an anomalous region is introduced into the model, making it more sensitive and responsive to such anomalies.
-This increased sensitivity is attributed to the improved performance of OmniAD.
-We conducted experiments to validate this hypothesis and demonstrate the efficacy of $A^{i}$.
 Figure 6: Student attention module.
@@
 Figure 7: ERF of OmniAD with and without the student attention module. (a) with $A^{I}$. (b) w/o $A^{I}$.
-IV ToyCity Dataset
+4 ToyCity Dataset
 The ToyCity dataset addresses the need for a benchmark dataset consisting of real images of multiple objects for Scene AD.
@@
 ToyCity features scenes of a toy city with buildings, signs, roads, trees, and miscellaneous items.
-The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure8.
-Each setup also contains different anomalies, as shown in Figure8.
-Additoinally, each anomaly image has at most one anomaly.
+The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure  8.
+Each setup also contains different anomalies, as shown in Figure  8.
+Additionally, each anomaly image has at most one anomaly but each setup has varying anomalies that were added and removed.
 The dataset includes the manually annotated ground truth for anomaly segmentation.
-Table I provides an overview of the number of images released for ToyCity.
-The dataset’s v1 is the base variant of all four scenes without any augmentations, v2 includes Quer-Aligned Novel Views (QANV) images rendered using Nerfacto, v3 incorporates augmentations using Interpolated Non-Anomalous Views (INV), and v4 combines the data of both v2 and v3, these variations are presented in Table II.
+Table 1 provides an overview of the number of images released for ToyCity.
+The dataset’s v1 is the base variant of all four scenes without any augmentations, v2 includes Quer-Aligned Novel Views (QANV) images rendered using Nerfacto, v3 incorporates augmentations using Interpolated Non-Anomalous Views (INV), and v4 combines the data of both v2 and v3, these variations are presented in Table 2.
 ToyCity thus offers a new benchmarking challenge to current state-of-the-art anomaly detection models.
@@
 A zoomed comparison is also presented that shows the anomaly highlighted with a red outline.
-TABLE I: Number of non-anomalous and anomalous images in each of the ToyCity dataset scenes, and the number of NVS images generated from each of the proposed methods QANV and IBV
+Table 1: Number of non-anomalous and anomalous images in each of the ToyCity dataset scenes, and the number of NVS images generated from each of the proposed methods QANV and IBV
 Name Non-Anomalous Anomalous QANV IBV
@@
 ToyCity 4 314 314 314 3757
-V Experiments
-We evaluate several AD methods on ToyCity and MAD to test the localization ability of each method.
+5 Experiments
+We evaluate OmniAD’s performance, its ablated variants, and several AD methods on ToyCity and MAD to test the localization ability of each method.
 We could only utilize a subset of the MAD dataset that consisted of real images compatible with COLMAP for Structure from Motion (SfM).
@@
 Finally a qualitative comparison between OmniAD and RD for anomaly segmentation is presented.
-V-A Metrics
+5.1 Metrics
 To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$ Score as described by Equation 1.
@@
 This balance is critical for datasets such as ToyCity and MAD, where anomaly regions typically constitute a small fraction of the total image area.
-Examples of the relatively small anomaly regions within ToyCity can be seen in Figure8, underscoring the necessity for such a balanced evaluation metric.
+Examples of the relatively small anomaly regions within ToyCity can be seen in Figure  8, underscoring the necessity for such a balanced evaluation metric.
 The imbalance of anomaly regions renders other metrics like accuracy or the area under the receiver operating characteristic curve (AUROC) as misleading.
@@
 Thus we provide both the $F_{1}$ score and AUROC.
-V-B Implementation Details
+5.2 Implementation Details
 To allow for fair comparison between all anomaly detection methods we use the anomalib [39] library implementation of each anomaly detection method.
@@
 Other hyper-parameters were left as default values provided by anomalib for each anomaly detection method.
-V-C Ablation Studies
-V-C1 OmniAD architecture
+5.3 Ablation Studies
+We conducted ablation studies to validate both the OmniAD architecture and the proposed NVS augmentation strategy.
+5.3.1 OmniAD architecture
 OmniAD has two major changes from RD.
@@
 To test if our changes are significant we test OmniAD with and without ResNeXt (OmniAD-R) and without $A^{i}$ (OmniAD-$A^{i}$) against RD.
-As shown in the $F_{1}$ scores on ToyCity Tables III and MAD Tables V we can see a consistent performance increase when $A^{i}$ is used.
-V-C2 NVS Augmentation
+As shown in the $F_{1}$ scores on ToyCity Tables 3 and 4 and MAD Tables 7 we can see a consistent performance increase when $A^{i}$ is used.
+5.3.2 NVS Augmentation
 We test the efficacy of our NVS augmentation strategies by evaluating our metrics with and without each of the augmentations.
-Our hypothesis that generating novel views of the our non-anomalous scene similar to the view point of the query scene will make it easier for OmniAD to identify anomalies is verified through four different variations listed in Table II.
-V-D Quantitative Evaluation
+Our hypothesis that generating novel views of the our non-anomalous scene similar to the view point of the query scene will make it easier for OmniAD to identify anomalies is verified through four different variations listed in Table 2.
+5.4 Quantitative Evaluation
 In our assessment of diverse AD methods on the ToyCity datasets, we observe a consistent pattern of performance increase OmniAD, OmniAD without ResNext (OmniAD w/o R), and RD when a novel view selection strategy is employed for $F_{1}$ scores on ToyCity and MAD.
-A similar trend can also be seen for AUROC on both ToyCity and MAD Tables IV,VI.
+A similar trend can also be seen for AUROC on both ToyCity and MAD Tables 5,6,8.
 Additional experiments on different scenes is provided in the supplemental material.
@@
 As mentioned metrics such as pixel-level AUROC can often show very high numbers for datasets such as ToyCity and MAD due to the low number of pixels that may be an anomaly.
-This can be seen for both ToyCity Table IV and MAD Table VI.
+This can be seen for both ToyCity Table 5,6 and MAD Table 8.
 Most AD methods will achieve an AUROC very close to 1 while still offering low $F_{1}$ scores.
-TABLE II: Four variations are tested for each of the scenes present in ToyCity.
+Table 2: Four variations are tested for each of the scenes present in ToyCity.
 Variation v1 is un-augmented, v2 is with the Query-Aligned Views, v3 is with the interpolated Non-Anomalous views, and v4 is the combination of Query-Aligned Baseline views and Interpolated Non-Anomalous views.
@@
 v4
-V-E Qualitative Evaluation
-Figure9 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, RD, which utilizes a ResNeXt backbone.
+5.5 Qualitative Evaluation
+Figure  9 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, RD, which utilizes a ResNeXt backbone.
 A threshold was selected for each method that maximizes their pixel-level $F_{1}$ score.
@@
 Figure 10: Additional segmentations and anomaly scorces from OmniAD.
-TABLE III: Comparison of Pixel $F_{1}$ scores on ToyCity 1, 2, 3, and 4
-Method ToyCity 1 ToyCity 2 ToyCity 3 ToyCity 4
-No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
-OmniAD 0.432 0.450 0.497 0.496 0.348 0.599 0.491 0.577 0.281 0.384 0.319 0.366 0.379 0.541 0.390 0.523
-OmniAD-R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
-OmniAD-$A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
-RD 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
-CFLOW 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
-FastFlow 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
-CFA 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
-DRÆM 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
-TABLE IV: Comparison of Pixel AUROC scores on ToyCity 1, 2, 3, and 4
-Method ToyCity 1 ToyCity 2 ToyCity 3 ToyCity 4
-No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
-OmniAD 0.990 0.996 0.999 0.999 0.991 0.998 0.999 0.999 0.979 0.993 0.997 0.999 0.982 0.993 0.995 0.998
-OmniAD-R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
-OmniAD-$A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
-RD 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
-CFLOW 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
-FastFlow 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
-CFA 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
-DRÆM 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
-TABLE V: Comparison of Pixel $F_{1}$ scores on MAD-Real [8] of Puppy, Pig, and Elephant
+Table 3: Comparison of Pixel $F_{1}$ scores on ToyCity 1 and 2
+Method ToyCity 1 ToyCity 2
+No Aug QANV INV Both No Aug QANV INV Both
+OmniAD 0.432 0.450 0.497 0.496 0.348 0.599 0.491 0.577
+OmniAD-R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503
+OmniAD-$A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357
+RD 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390
+CFLOW 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082
+FastFlow 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128
+CFA 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081
+DRÆM 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039
+Table 4: Comparison of Pixel $F_{1}$ scores on ToyCity 3 and 4
+Method ToyCity 3 ToyCity 4
+No Aug QANV INV Both No Aug QANV INV Both
+OmniAD 0.281 0.384 0.319 0.366 0.379 0.541 0.390 0.523
+OmniAD-R 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
+OmniAD-$A^{i}$ 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
+RD 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
+CFLOW 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
+FastFlow 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
+CFA 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
+DRÆM 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
+Table 5: Comparison of Pixel AUROC scores on ToyCity 1 and 2
+Method ToyCity 1 ToyCity 2
+No Aug QANV INV Both No Aug QANV INV Both
+OmniAD 0.990 0.996 0.999 0.999 0.991 0.998 0.999 0.999
+OmniAD-R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999
+OmniAD-$A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997
+RD 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998
+CFLOW 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868
+FastFlow 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952
+CFA 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984
+DRÆM 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725
+Table 6: Comparison of Pixel AUROC scores on ToyCity 3 and 4
+Method ToyCity 3 ToyCity 4
+No Aug QANV INV Both No Aug QANV INV Both
+OmniAD 0.979 0.993 0.997 0.999 0.982 0.993 0.995 0.998
+OmniAD-R 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
+OmniAD-$A^{i}$ 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
+RD 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
+CFLOW 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
+FastFlow 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
+CFA 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
+DRÆM 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
+Table 7: Comparison of Pixel $F_{1}$ scores on MAD-Real [13] of Puppy, Pig, and Elephant
 Method Puppy Pig Elephant
@@
 DRÆM 0.042 0.016 0.022 0.018 0.040 0.041 0.072 0.110 0.020 0.041 0.037 0.028
-TABLE VI: Comparison of Pixel AUROC scores on MAD-Real [8] of Puppy, Pig, and Elephant
+Table 8: Comparison of Pixel AUROC scores on MAD-Real [13] of Puppy, Pig, and Elephant
 Method Puppy Pig Elephant
@@
 DRÆM 0.819 0.359 0.514 0.524 0.720 0.571 0.684 0.796 0.601 0.496 0.559 0.519
-VI Conclusion
+6 Limitations and Future Work
+OmniAD and the proposed ToyCity dataset address unsupervised Scene Anomaly Detection in multi-view scenarios, yet several limitations remain.
+First, each ToyCity image contains at most one anomaly at a time, even though multiple anomalies were added and removed from each scene.
+In contrast, real-world settings such as disaster assessments or industrial inspections may involve multiple anomalies within a single capture.
+Note that ToyCity does include partial occlusions, and OmniAD itself does not strictly limit detection to a single anomaly region.
+Second, although ToyCity was captured with real cameras at varying angles and lighting conditions, its tabletop-scale scenes do not fully encompass the complexities of large scale infrastructure.
+Furthermore, our data augmentation assumes sufficient coverage for camera localization.
+In highly variable environments or when dealing with reflective surfaces, generating high-quality synthetic views can be challenging without accurate camera localization.
+Nonetheless, as shown in our experiments, OmniAD’s performance does not hinge solely on these augmented views.
+In future work, we plan to:
+  • 1.
+    Extend ToyCity to include multiple anomalies in a single image and introduce dynamic elements such as moving vehicles or pedestrians.
+  • 2.
+    Validate OmniAD on larger real-world datasets (e.g., complete bridge or ship inspections) to further assess its robustness.
+  • 3.
+    Investigate faster 3D reconstruction, localization, and noval view synthesis methods.
+  • 4.
+    Identifying the underlying failure modes in GAN/autoencoder based AD methods under significant viewpoint variation remains an open research question, which we leave for future work.
+7 Conclusion
 Our study introduced and addressed a specific anomaly detection (AD) problem we term Scene AD.
-Scene AD deals with pixel-level AD between a non-anomalous scene and an anomalous scene, where a misalignment between non-anomalous and anomalous views is expected, as is often the case in inspections of uncontrolled environments.
+Scene AD deals with pixel-level AD between a non-anomalous scene and an anomalous scene, where a misalignment between non-anomalous and anomalous views is expected, as is often the case in applications in the built environment.
 To challenge and measure the effectiveness of state-of-the-art anomaly detection for Scene AD, we introduced the ToyCity dataset.
 This new benchmark focuses on viewpoint variance in scenes of multiple objects and serves as a testbed for our new method OmniAD.
-We demonstrated through extensive comparisons that OmniAD outperforms other AD methods when applied to the ToyCity dataset, establishing a new state-of-the-art for this task.
+We demonstrated through extensive comparisons that OmniAD outperforms other AD methods when applied to the ToyCity dataset as well as the object centric established MAD-Real, establishing a new state-of-the-art for this task.
 Particularly, we found that the incorporation of a student attention module increases the efficacy of reverse distillation for pixel-level AD by expanding the effective receptive filed (ERF).
-Our experiments also demonstrated the benefits of novel view synthesis and camera localization for dataset augmentation in enhancing the robustness and accuracy of OmniAD.
+Our experiments also demonstrated the benefits of novel view synthesis and camera localization for dataset augmentation in enhancing the robustness and accuracy of OmniAD allowing for an increase in pixel level $F_{1}$ of up to 40% in many scenes.
 We hope that our work will catalyze further advancements, spurring the creation of advanced anomaly detection techniques capable of managing the complexities associated with varying viewpoints in 3D scenes.
+8 Acknowledgments
+The authors acknowledge partial financial support from the Texas Department of Transportation (Grant No.
+0-7181), and Department of Defense (Project No.
+G0511607).
+The contents of this paper reflect the views of the authors, who are responsible for the facts and the accuracy of the data presented herein.
+The contents do not necessarily reflect the official views or policies of the sponsors.
+The authors acknowledge the use of the Carya Cluster and the advanced support from the Research Computing Data Core at the University of Houston to carry out the research presented here.

### v3

@@
 Abstract
-Visual anomaly detection in the built environment can be a valuable tool for applications such as infrastructure assessment, construction monitoring, security surveillance, and urban planning.
-Anomaly detection approaches are typically unsupervised and work by detecting deviations from an expected state where no assumptions are made exact type of deviation.
-Unsupervised pixel-level anomaly detection methods have been developed to successfully recognize and segment anomalies; however, existing techniques are designed for industrial settings with a fixed camera position.
-In the built environment, images may be periodically collected by a moving camera manually or by aerial or ground vehicles and utilized for decision making.
-The camera pose between successive collections may vary widely voiding a fundamental assumption in existing anomaly detection approaches.
-To address this gap, we introduce the problem of Scene Anomaly Detection (Scene AD), where the goal is to detect anomalies from two sets of images: one set without anomalies and one set that may or may not contain anomalies.
-No labeled semantic segmentation data are provided for training.
-We propose a novel network, OmniAD, to tackle Scene AD by refining the reverse distillation anomaly detection method, leading to a 40% improvement in pixel-level anomaly detection.
-Additionally, we introduce two new data augmentation strategies that leverage novel view synthesis and camera localization to enhance generalization.
-We evaluate our approach both qualitatively and quantitatively on a new dataset, ToyCity the first Scene AD dataset featuring multiple objects as well as on the established single object centric dataset, MAD.
-Our method demonstrates marked improvement over baseline approaches, paving the way for robust anomaly detection in scenes with real-world camera pose variations commonly observed in the built environment.
+The built environment, encompassing critical infrastructure such as bridges and buildings, requires diligent monitoring of unexpected anomalies or deviations from a normal state in captured imagery.
+Anomaly detection methods could aid in automating this task; however, deploying anomaly detection effectively in such environments presents significant challenges that have not been evaluated before.
+These challenges include camera viewpoints that vary, the presence of multiple objects within a scene, and the absence of labeled anomaly data for training.
+To address these comprehensively, we introduce and formalize Scene Anomaly Detection (Scene AD) as the task of unsupervised, pixel-wise anomaly localization under these specific real-world conditions.
+Evaluating progress in Scene AD required the development of ToyCity, the first multi-object, multi-view real-image dataset, for unsupervised anomaly detection.
+Our initial evaluations using ToyCity revealed that established anomaly detection baselines struggle to achieve robust pixel-level localization.
+To address this, two data augmentation strategies were created to generate additional synthetic images of non-anomalous regions to enhance generalizability.
+However, the addition of these synthetic images alone only provided minor improvements.
+Thus, OmniAD, a refinement of the Reverse Distillation methodology, was created to establish a stronger baseline.
+Our experiments demonstrate that OmniAD, when used with augmented views, yields a 64.33% increase in pixel-wise $F_{1}$ score over Reverse Distillation with no augmentation.
+Collectively, this work offers the Scene AD task definition, the ToyCity benchmark, the view synthesis augmentation approaches, and the OmniAD method.
 Project Page: https://drags99.github.io/OmniAD/
@@
 Anomaly Detection , Unsupervised Learning , Camera Localization , Neural Radiance Fields
+journal:
 […]
@@
 1 Introduction
-Figure 1: In Scene AD, we observe variations in camera positions between non-anomalous views (green), and query views which may contain anomalies (red).
+Detecting anomalies is critical for many scene assessment and analysis applications in the built environment, including infrastructure inspections[1, 2, 3], disaster response[4, 5], construction management [6] and planning [7].
+Anomalies can be described as a deviation (e.g., damage, structural changes, new objects) from an established non-anomalous reference.
+In many scene assessment and analysis applications, images are the primary sensing modality, as they are inexpensive to capture and provide the dense resolution required for visual evidence in documentation [8, 9].
+Anomaly detection (AD) approaches could be categorized as supervised or unsupervised.
+Supervised approaches rely on annotated data and work well when high-quality annotations are available but can be impractical in many real-world inspection scenarios where anomalies are rare or unknown in advance [10, 11, 12].
+Unsupervised AD methods, on the other hand, are often more practical as they don’t require labeled data of anomalies.
+Current benchmarks for unsupervised anomaly detection have largely been tailored for manufacturing contexts which consist of imagery of a single-object [13, 14, 12] or multiple objects [15, 16] with constrained views.
+While unsupervised image-based AD methods have found great success in manufacturing applications, they have not been explored for scene analysis applications in the built environment, where many manufacturing-specific assumptions no longer hold.
+Imagery in the built environment is typically collected from unconstrained viewpoints, often by handheld or UAV-mounted cameras [1, 17, 18, 19].
+As shown in Figure 1 view-points for images captured in the built environment vary in nature tremendously, unlike the fixed-camera setting typically used for unsupervised AD benchmarks [13, 12, 20, 15].
+Seminal anomaly detection approaches [21, 22, 23, 24, 25] discussed in more detail in the related works section have matured to the point of achieving very high quantitative metrics with fixed views, with state-of-the-art methods plateauing in performance.
+However, our studies showed that seminal works are unable to meet similar performance levels for multi-view multi-object scenes such as in Figure 1 (a).
+This performance disparity underscores the need to specifically define and address the problem of anomaly detection within these unconstrained visual environments.
+An effective scene anomaly detection method in unconstrained visual environments must operate (i) without labeled data, (ii) across multiple objects per image, and (iii) under random viewpoint variation.
+We formally refer to this general AD task as Scene Anomaly Detection (Scene AD) and illustrate it in Figure 2.
+Formally, Scene AD entails localizing anomalous regions given two unlabeled image sets: one depicting normal reference conditions, called the non-anomalous set, and one containing potential anomalies, if any, called the query set.
+No assumption is made for alignment between the non-anomalous and query set.
+Camera poses, if needed, should only be estimates from methods such as structure-from-motion or GPS measurements to mimic the noisy camera poses that can feasibly be obtained in an uncontrolled environment.
+To provide a benchmark for Scene AD, we introduce ToyCity a real-world dataset consisting of four multi-object scenes, with more than 1200 anomalous images per scene and over 20 varying anomalies across scenes.
+The diversity of anomalies and quantity of anomalous images in ToyCity match or exceed those of the most frequently used fixed view AD benchmarks [15, 26, 16, 27, 13, 20, 14].
+ToyCity also exceeds the real-image diversity of the Multi-Pose Anomaly Detection (MAD)[12] dataset, which contains 16 to 32 anomalous images per object.
+When evaluating baseline AD methods, that perform well on prior datasets, we found they exhibit very weak pixel-level localization on ToyCity.
+Among the baseline AD methods, Reverse Distillation (RD) [25] showed the best localization performance, however we found its localization performance to be inconsistent for practical use.
+We introduce OmniAD to validate the potential of unsupervised AD approaches for pixel-level Scene Anomaly Detection.
+OmniAD is a refinement of Reverse Distillation [25] that integrates student-attention modules.
+These modules are designed to expand the student decoder’s effective receptive field (ERF) [28], thereby improving its sensitivity to localized anomalies.
+Furthermore, to counteract the performance degradation caused by viewpoint variations in Scene AD, we explore two data augmentation strategies.
+Both strategies aim to generate novel non-anomalous views, using Neural Radiance Fields (NeRF) [29] and hierarchical localization (hloc) [30, 31], to assist AD methods in generalizing to varying views.
+Figure 1: (a) Typical images used in the built environment.
+(b) Images available for benchmarking unsupervised anomaly detection [13, 15, 16].
+The top row for (a) and (b) consist of non-anomalous images, while the bottom is anomalous.
+Figure 2: In Scene AD, we observe variations in camera positions between non-anomalous views (green) and query views that may contain anomalies (red).
 In contrast, traditional AD methods typically feature aligned views between the non-anomalous and anomalous images.
-For many applications in the built environment —ranging from infrastructure assessment and construction monitoring to security surveillance and urban planning—detecting visual anomalies in scenes captured over time is a critical task  [1, 2, 3, 4, 5].
-These anomalies may arise from structural changes, newly introduced objects, unexpected modifications in appearance, or more generally, anything that deviates from an established non-anomalous reference.
-However, images acquired at different instances, whether manually or autonomously, typically suffer from misalignment due to variations in camera pose.
-For example, bridge inspectors may assess changes to a structure following a disaster, self-driving car systems might detect modifications to traffic signs, and the U.S.
-Military places strong emphasis on monitoring infrastructure for anomalies critical to operational readiness [6, 7, 8, 9, 10, 11, 12].
-In all such cases, each new set of images is unaligned with previous captures, complicating anomaly identification.
-We define the task of Scene Anomaly Detection (Scene AD) as detecting anomalous regions in a query set of images, potentially containing no anomalies at all, against a non-anomalous reference set.
-Labeled examples of what qualifies as an anomaly are unavailable, rendering supervised segmentation methods impractical.
-Additionally anomalies are often too infrequent or unpredictable to compile sufficient labeled training data for every potential anomaly type, rendering fully supervised models infeasible.
-As a result, unsupervised anomaly detection (AD) approaches are a feasible alternative.
-However, these AD methods were originally crafted for controlled industrial scenarios which commonly assume fixed camera views and struggle with the viewpoint variations of real-world Scene AD.
-Two main challenges emerge.
-First, there is no suitable multi-object, multi-view benchmark for Scene AD: most existing AD datasets capture a single view or focus on a single object [13].
-Second, any robust Scene AD solution must tolerate significant camera misalignment and still detect newly introduced or changed regions.
-To address these gaps, we propose OmniAD, a refinement of the Reverse Distillation (RD) framework [14].
-RD detects anomalies by comparing a student decoder’s reconstruction with a teacher encoder’s embedding.
-We hypothesize that broadening the student decoder’s effective receptive field (ERF) [15] will improve sensitivity to anomalous regions.
-Accordingly, we incorporate student attention modules to expand the ERF and enhance pixel-level anomaly detection.
-Moreover, to manage viewpoint differences in real scenes, we introduce two data augmentation strategies leveraging Neural Radiance Fields (NeRF) [16] and hierarchical localization (hloc) [17, 18].
-These techniques generate novel, more closely aligned views of the query scene, aiding unsupervised methods—such as OmniAD—in detecting anomalies without labeled examples.
-Since no comprehensive benchmark exists for this multi-object, multi-view setting, we introduce the ToyCity dataset.
-ToyCity consists of scenes with multiple objects, varied camera viewpoints, and different anomalies.
-We evaluate OmniAD and several baseline methods on ToyCity as well as on the single-object-centric Multi-pose Anomaly Detection dataset (MAD) [13], demonstrating notable performance gains.
 Our main contributions are:
   • 1.
-    Scene AD Task: Formal introduction of Scene Anomaly Detection, highlighting the challenge of detecting anomalies in unaligned multi-view imagery.
+    Scene AD Task: Formal introduction of Scene Anomaly Detection, highlighting the challenge of detecting anomalies in unaligned multi-view imagery consisting of multiple objects.
   • 2.
-    ToyCity Dataset: A new multi-object, multi-view dataset designed for benchmarking Scene AD methods in realistic settings.
+    ToyCity Dataset: A new multi-object multi-view dataset of real images designed for benchmarking methods for Scene AD.
   • 3.
-    OmniAD Architecture: A refined Reverse Distillation approach with a proposed Student Attention Module to expand the ERF, achieving state-of-the-art performance on ToyCity and MAD [13] and validated with ablation studies.
+    Baselines: We establish the first quantitative metrics of seminal AD methods for Scene AD using ToyCity.
   • 4.
-    NeRF-based Data Augmentation Methods: Two novel techniques using NeRF and hierarchical localization for improved view-invariance, yielding a 40% increase in anomaly localization accuracy.
+    NeRF based Data Augmentation Methods: Two augmentation strategies for increased view-generalization in Scene AD.
+  • 5.
+    OmniAD: A refinement of Reverse Distillation [25], which when combined with augmentation strategies leads to a 64.33% improvement in pixel-wise $F_{1}$ score for Scene AD compared to Reverse Distillation with no augmentation.
 The rest of the paper is organized as follows: Section 2 reviews related work on anomaly detection, neural radiance fields, and camera localization.
@@
 Section 6 discusses the limitations of our dataset and method.
-Finally, Section 7 concludes the paper.
+Finally, Section 7 summarizes and concludes this paper.
 2 Related Work
-Anomaly Detection.
-AD algorithms typically harness a collection of non-anomalous images , to establish a normative distribution.
+2.1 Unsupervised Image-Based Anomaly Detection
+Anomaly detection (AD) algorithms typically harness a collection of non-anomalous images, to establish a normative distribution.
 AD is a critical function in industrial visual inspection for identifying defects [10] and is also employed in healthcare for the detection of irregularities in medical images [32].
-Prominent AD methods include CFLOW-AD [21], FastFlow [22], CFA [23], DRÆM [24], and Reverse Distillation (RD) [14].
+Unsupervised AD methods have demonstrated considerable success, particularly for fixed-view single-object anomaly detection.
+This domain has become mature over time to achieve a high-performance plateau for pixel-wise anomaly detection.
+New methods [33, 34, 35, 36] frequently benchmark against a core group of foundational techniques [21, 22, 23, 24, 25] due to the high-performance plateau achieved.
+When introducing novel tasks such as Scene AD, we must select robust baseline methods that are demonstrably reproducible and widely accepted for comparison to create a meaningful benchmark.
+Accordingly, our analysis centers on prominent methods such as CFLOW-AD [21], FastFlow [22], Coupled-hypersphere-based Feature Adaptation (CFA) [23], discriminatively trained reconstruction anomaly embedding model (DRÆM) [24], and Reverse Distillation (RD) [25].
+These specific methods cover the two primary categories unsupervised image-based AD methods fall into, feature embedding based and reconstruction based [12, 34].
+Furthermore, this strategic selection aims to establish a durable and widely understood performance benchmark.
+These methods offer strong guarantees of reproducibility, stemming from their continuous validation and the availability of robust, often multiple, implementations [37].
+2.2 Feature Embedding Based Unsupervised Anomaly Detection
 Both CFLOW-AD and FastFlow utilize a flow framework [38] that models the distribution of normal image patches to detect anomalous patterns.
@@
 CFA fine-tunes the pretrained CNN features through metric learning, which is based on a coupled hypersphere approach, clustering normal features tightly within a hypersphere’s boundary.
+2.3 Reconstruction Based Unsupervised Anomaly Detection
 DRÆM operates with two main networks: a reconstructive network and a discriminative network.
@@
 The discriminative network uses both the reconstructed normal image and the synthetic anomalous image to predict segmentation for the regions where the synthetic anomaly was introduced.
-RD adopts a knowledge distillation strategy [26, 27], where a student decoder network learns the feature space representation from a teacher encoder network at various feature map scales.
-Unlike traditional approaches that feed the input image into both the student and the teacher networks, RD takes an autoencoder-like approach where the teacher encodes the input image, and the student decodes a fused multi-scale feature map output by the teacher.
-This feature map is a compressed representation of the teacher’s features across multiple scales.
-Predominant datasets for benchmarking these methods, like MVTecAD [19] and VisA [28], typically employ a fixed overhead camera perspective, resulting in a uniform view of anomalies.
-Our research demonstrates that significantly different views can lead to substantial degradation in the pixel-level localization of anomalies.
-To combat these limitations, our method leverages a novel view synthesis method and camera localization techniques to generate corresponding non-anomalous images, thus facilitating improved anomaly detection.
-While there are existing methods for 3D anomaly detection, they generally require 3D data that is aligned with a reference model.
-Neural Radiance Field NeRFs, as introduced by Mildenhall et al.
-[16], represent a significant advancement in the field of view synthesis.
+Reverse Distillation adopts a knowledge distillation strategy [39, 40], where a student decoder network learns the feature space representation from a teacher encoder network at various feature map scales.
+Unlike traditional approaches that feed the input image into both the student and the teacher networks, Reverse Distillation takes an autoencoder-like approach where the teacher encodes the input image, and the student decodes a fused multiscale feature map output by the teacher.
+Deviations between the student and teachers’ features indicate the occurrence of an anomaly.
+2.4 Multi-View Anomaly Detection
+While the aforementioned unsupervised AD methods demonstrate strong performance in controlled environments, a significant limitation inherent in many of these 2D approaches arises when deployed in multi-view settings [12].
+Their effectiveness often depends critically on feature representations or learned normality models (such as distributions, reconstruction targets, or feature manifolds) that are derived from training data [10, 11].
+As a result, these methods frequently struggle when presented with the significant viewpoint variations from those seen in the training data that is typical of real-world data acquisition, like images captured using handheld devices or UAVs during infrastructure assessments [1, 19].
+Additionally, realistic scenes often contain multiple objects simultaneously, a scenario largely distinct from typical single-object AD benchmarks.
+Consequently, the specific task of unsupervised anomaly detection within environments characterized by both multiple objects per scene and random, unconstrained multi-view captures has remained largely unexplored.
+This is partly due to the lack of suitable benchmark datasets designed to address these combined complexities, as the manual effort in creating such a dataset with real images is tremendous in comparison to labeling the occurrence of an anomaly on a single object.
+2.5 Neural Radiance Field
+Neural Radiance Fields (NeRFs) [29] represent a significant advancement in the field of view synthesis.
 These fields use a collection of images with known camera poses to synthesize new viewpoints of a scene.
@@
 Subsequent innovations like Nerfacto [41] and Instant-ngp [42] have further refined NeRFs, offering rapid training and real-time rendering capabilities.
-In the context of anomaly detection, we leverage NeRFs to synthesize non-anomalous images from diverse viewpoints, significantly improving the detection process.
-Hierarchical Localization The process of hierarchical localization (hloc) [31, 18] has been developed to accurately determine a camera’s position and orientation in space, a task known as 6-Degree-of-Freedom (DoF) localization [31, 18, 32].
+In the context of this work, we leverage NeRFs to synthesize non-anomalous images from diverse viewpoints.
+2.6 Hierarchical Localization
+The process of hierarchical localization (hloc) [43, 31] has been developed to accurately determine a camera’s position and orientation in space, a task known as 6-Degree-of-Freedom localization [43, 31, 44].
 This process begins with the extraction of local features from a query image, utilizing methods such as SuperPoint [30], Sift [45], or DISK [46].
 A 3D SfM model serves as a reference for these local features, against which the query image is matched using global descriptors like VLAD [47].
-The Perspective-n-Point algorithm, in conjunction with RANSAC, then utilizes these correspondences to estimate the camera’s pose.
+The Perspective-n-Point algorithm, with RANSAC, then utilizes these correspondences to estimate the camera’s pose.
 This pose is critical in enabling NeRFs to generate non-anomalous images that correspond to those containing anomalies, thus in theory providing a more robust dataset for anomaly detection algorithms.
-However as we shall show an improved AD method is needed to fully utilize these anomaly aligned images.
+However, as we shall show, an improved AD method is needed to fully utilize these anomaly aligned images.
 3 Method
-An overview of our AD methodology for Scene AD can be seen in Figure  2.
+An overview of our AD methodology for Scene AD can be seen in Figure 3.
 Scene AD consists of non-anomalous views and query views.
 We generate an SfM model using non-anomalous views to acquire camera poses.
-Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation as shown in Figure  3.
+Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation, as shown in Figure 4.
 For the first strategy, we interpolate between the non-anomalous camera poses.
 The second strategy for generating useful camera views for training is to utilize hloc to acquire the camera poses of query views.
-We then use NVS with the new camera poses to augment our non-anomalous views for training our proposed method OmniAD.
-The following subsections describes each of these steps in more detail.
+We then use a NeRF for novel view synthesis (NVS) with the new camera poses to augment our non-anomalous views for training our proposed method OmniAD.
+The following subsections describe each of these steps in more detail.
 Figure 3: Overview of our methodology is depicted here.
@@
 Leveraging these non-anomalous camera poses, we design a trajectory that connects them, allowing for interpolation.
-Subsequently, we apply a NVS technique to generate images that present unique perspectives along the interpolated path.
-The anomaly dataset can also be used to determine more novel views.
+Subsequently, we utilize a NeRF for Novel View Synthesis (NVS) to generate images that present unique perspectives along the interpolated path.
+Query images can also be used to determine more novel views.
 3.1 Novel View Selection Strategy
-To generate new non-anomalous camera poses for NVS, we propose and evaluate two different approaches, namely Interpolated Non-anomalous Views (INV) and Query-aligned non-anomalous views (QANV).
-3.1.1 Interpolated Non-anomalous Views
-Figure 4: The provided sequence illustrates camera interpolation, where the pose transitions from the starting image to the ending image, are used to synthesize intermediate views.
-These interpolated views, not present in the original dataset, bridge the gap between the start and end poses, introducing new perspectives to the dataset.
-Given a set of non-anomalous images an associated camera poses $T$ can be calculated using SfM.
+A core challenge in Scene AD is the viewpoint variation between the non-anomalous reference images and the query images.
+Anomaly detection models trained only on the original non-anomalous views may struggle when presented with a query image from a significantly different perspective.
+To enhance the model’s robustness to these variations, a wider variety of non-anomalous viewpoints are required for training, ideally including views similar to those encountered in the query set.
+To generate these new non-anomalous camera poses for novel view synthesis (NVS), we propose and evaluate two different approaches:
+  • 1.
+    Interpolated Non-anomalous Views (INV): The goal of this strategy is to densify the coverage of the non-anomalous viewpoint space.
+By generating views between the existing captured non-anomalous poses, we train the model to handle gradual shifts in perspective, improving its general robustness to unseen viewpoints near the original capture trajectory.
+  • 2.
+    Query-Aligned Non-anomalous Views (QANV): This strategy directly aims to address the misalignment with potentially anomalous query images.
+By estimating the camera pose of a query image and then synthesizing a non-anomalous view from that approximate camera pose, we provide the model with similar normal/potentially anomalous pairs (in terms of viewpoint).
+Intuitively, this allows the model to better isolate deviations caused by anomalies rather than viewpoint changes.
+The following subsections detail how these camera poses are generated and used with NVS.
+3.1.1 Interpolated Non-anomalous Views (INV)
+Given a set of non-anomalous images, an associated camera poses $T$ can be calculated using SfM.
 The calculated $T$ can be interconnected to form a trajectory.
@@
 Specifically, for our experiments, we interpolate 12 intermediate camera poses between consecutive original poses in $T$, thereby enhancing the dataset with interpolated views that increase the robustness of our AD model to various camera angles.
-An example of the camera interpolation for ToyCity can be seen in Figure  4.
-3.1.2 Query-Aligned Non-anomalous Views
-Here we propose a method to generate images of the baseline scene, but using poses from the anomaly scene.
-Our method employs hloc to accurately estimate the camera pose for a query anomaly image.
-This necessitates an accurate reconstruction of the baseline scene that is provided by the baseline SfM model.
-Hloc provides a variety of options for local feature detection, matching, and image similarity measurement.
+An example of the camera interpolation for ToyCity can be seen in Figure  5.
+Figure 5: The provided sequence illustrates camera interpolation, where the pose transitions from the starting image to the ending image, are used to synthesize intermediate views.
+These interpolated views, not present in the original dataset, bridge the gap between the start and end poses, introducing new perspectives to the dataset.
+3.1.2 Query-Aligned Non-anomalous Views (QANV)
+Here, we propose a method to generate additional camera poses relative to the non-anomalous scene by localizing images from the query set.
+Our method employs hloc to estimate the camera pose for a query anomaly image.
+This necessitates an accurate reconstruction of the non-anomalous scene that is provided by the non-anomalous SfM model.
+Hloc provides various options for local feature detection, matching, and image similarity measurement.
 We opted for a combination of SuperPoint trained on Aachen and SuperGlue [30, 48], as it demonstrated effective performance on our datasets on average.
@@
 Continual advancements in this field, like the advent of 3D Gaussian Splattering[49], present opportunities for further enhancing rendering time and quality.
-Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure  3.
-Example camera interpolated images can be seen in Figure  3, and example images using QANV can be seen in Figure  3
+Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure 4.
+Example camera interpolated images can be seen in Figure 4, and example images using QANV can be seen in Figure 4
 3.3 OmniAD
 Our goal in Scene AD is to handle significant camera misalignment while detecting anomalies at a pixel level.
-Although Reverse Distillation (RD) [14] provides a robust teacher student mechanism for fixed view scenarios, our preliminary tests revealed its limitations in multi-view scenarios.
+Although Reverse Distillation[25] provides a robust teacher student mechanism for fixed view scenarios, our preliminary tests revealed its limitations in multi-view scenarios.
+As such, OmniAD is proposed to refine Reverse Distillation to improve its generalizability in Scene AD, an overview of OmniAD is shown in 6.
+Figure 6: OmniAD’s architecture mirrors RD’s, featuring a frozen teacher encoder that channels features through a bottleneck module into a student decoder.
+We enhance this setup by integrating student attention modules $A^{1}$, $A^{2}$, and $A^{3}$.
 Backbone Exploration.
 We first explored different backbones to replace RD’s default encoder, focusing on architectures that could accommodate the diverse viewpoints.
-Many alternatives struggled to converge or offered minimal gains, but ResNeXt [38] consistently yielded better multi-scale representations and improved performance compared to simpler encoders.
+Many alternatives struggled to converge or offered minimal gains, but ResNeXt[50] consistently yielded better multiscale representations and improved performance compared to simpler encoders.
 In principle, any sufficiently expressive architecture including fully attention-based backbones could be substituted for the encoder.
@@
 Receptive Field Constraints.
-Further analysis indicated that RD’s student decoder had a relatively small effective receptive field (ERF) [15], causing it to overlook anomalies.
+Further analysis indicated that Reverse Distillation’s student decoder had a relatively small effective receptive field (ERF) [28], causing it to overlook anomalies.
 This shortcoming proved critical in multi-view settings, where anomalies could appear in various positions or at different scales.
@@
 Let $\mathcal{I}$ denote the image domain (e.g., $H\times W$ pixels), and let $x\in\mathbb{R}^{C\times H\times W}$ be an input image.
-Suppose a teacher network $T$ extracts multi-scale features
-$T(x)\;=\;\bigl{(}T^{1}(x),T^{2}(x),\dots,T^{L}(x)\bigr{)},$
+Suppose a teacher network $T$ extracts multiscale features,
+$T(x)\;=\;\bigl{(}T^{1}(x),T^{2}(x),\dots,T^{L}(x)\bigr{)},$ (1)
 where $T^{l}(x)\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the $l$-th feature map.
 A student network $S$ decodes these teacher features to produce a reconstruction $\hat{x}=(S\circ T)(x)$.
-In Reverse Distillation [14], any mismatch between $x$ and $\hat{x}$ can indicate an anomaly, since $S$ is trained on non-anomalous (normal) data alone.
+In RD [25], any mismatch between $x$ and $\hat{x}$ can indicate an anomaly, since $S$ is trained on non-anomalous data alone $S$ is unable to reconstruct anomalous features from a teacher.
 Effective Receptive Field (ERF). For each output location $(i,j)$ in $\hat{x}$, define its effective receptive field as
-$\mathrm{ERF}_{ij}(S\circ T)\;=\;\bigl{\{}\,(u,v)\in\mathcal{I}\;\big{|}\;\tfrac{\partial\hat{x}_{ij}}{\partial x_{uv}}\neq 0\bigr{\}}.$
+$\mathrm{ERF}_{ij}(S\circ T)\;=\;\bigl{\{}\,(u,v)\in\mathcal{I}\;\big{|}\;\tfrac{\partial\hat{x}_{ij}}{\partial x_{uv}}\neq 0\bigr{\}}.$ (2)
 In standard RD, convolutions and limited skip connections yield a relatively localized ERF.
@@
 By adding student attention modules (Fig.
-6), the new student $S^{\prime}$ broadens $\mathrm{ERF}_{ij}(S^{\prime}\circ T)$:
-$\bigl{|}\mathrm{ERF}_{ij}(S^{\prime}\circ T)\bigr{|}\;\;\geq\;\;\bigl{|}\mathrm{ERF}_{ij}(S\circ T)\bigr{|}.$
+7), the new student $S^{\prime}$ broadens $\mathrm{ERF}_{ij}(S^{\prime}\circ T)$,
+$\bigl{|}\mathrm{ERF}_{ij}(S^{\prime}\circ T)\bigr{|}\;\;\geq\;\;\bigl{|}\mathrm{ERF}_{ij}(S\circ T)\bigr{|}.$ (3)
 Consequently, for any anomalous pixel $(u,v)\in A$, there are more $(i,j)$ in the reconstructed output $\hat{x}$ for which $(u,v)\in\mathrm{ERF}_{ij}(S^{\prime}\circ T)$.
 In other words, the anomaly influences a broader swath of the reconstruction, resulting in larger or more easily detectable errors.
-Reverse Distillation Anomaly Score.
-As in conventional RD, we define an anomaly score per pixel $(i,j)$ by measuring discrepancy between $\hat{x}_{ij}$ and $x_{ij}$.
+Effect of ERF on Anomaly Score.
+As in conventional RD, we define an anomaly score per pixel $(i,j)$ by measuring the discrepancy between $\hat{x}_{ij}$ and $x_{ij}$.
 When the ERF is small, an anomaly’s influence might be confined or even “explained away” by the network averaging out the region, yielding low detection sensitivity.
 With an expanded ERF, the anomaly is more likely to cause a significant mismatch for at least one output location $(i,j)$, thereby increasing detection rates.
-Attention Modules for the Student Decoder.
+Expansion of ERF with Student Attention Modules.
 To address the ERF limitation, we introduced additional attention layers, denoted $A^{i}$, after each decoder block.
-These modules expand the ERF by allowing the network to aggregate information from spatially distant regions, thus capturing subtle or pose-dependent anomalies.
-Figure 6 shows the basic layout of each attention module, incorporating interleaved 3x3 convolutions and self-attention blocks for both local and global feature integration.
-Figure 5: OmniAD’s architecture mirrors RD’s, featuring a frozen teacher encoder that channels features through a bottleneck module into a student decoder.
-We enhance this setup by integrating student attention modules $A^{1}$, $A^{2}$, and $A^{3}$, bolstering the architecture’s proficiency in utilizing Query-aligned non-anomalous views and Interpolated Baseline Views for improved effectiveness.
-Figure 6: Student attention module.
+These modules expand the ERF, as shown in Figure 8, by allowing the network to aggregate information from spatially distant regions, thus capturing subtle or pose-dependent anomalies.
+Figure 7 shows the basic layout of each attention module and relevant parameters are shown in Table 1.
+Table 1: Student attention module parameters and layers: $A^{3}$: $C_{in}=256,H_{in}=64,C_{out}=256,H_{out}=64$.
+$A^{2}$: $C_{in}=512,H_{in}=32,C_{out}=512,H_{out}=32$.
+$A^{1}$: $C_{in}=1024,H_{in}=16,C_{out}=1024,H_{out}=16$.
+Batch Normalization (BN), 2D Max Pooling (MaxPool2D), 2D Convolution (Conv2D)
+Stage Layer Type Details Output Shape
+Input - Input Channels: $C_{in}$, Spatial Size: $H_{in}\times H_{in}$ $B\times C_{in}\times H_{in}\times H_{in}$
+Encoder Block 1 Conv2D kernel=3, stride=1, padding=1, $C_{in}\to 64$ $B\times 64\times H_{in}\times H_{in}$
+Self Attention channels=64 $B\times 64\times H_{in}\times H_{in}$
+MaxPool2D kernel=2, stride=2 $B\times 64\times H_{in}/2\times H_{in}/2$
+Encoder Block 2 Conv2D kernel=3, stride=1, padding=1, $64\to 128$ $B\times 128\times H_{in}/2\times H_{in}/2$
+Self Attention channels=128 $B\times 128\times H_{in}/2\times H_{in}/2$
+MaxPool2D kernel=2, stride=2 $B\times 128\times H_{in}/4\times H_{in}/4$
+Bottleneck Conv2D, BN, ReLU kernel=3, stride=1, padding=1, $128\to 256$ $B\times 256\times H_{in}/4\times H_{in}/4$
+Self Attention channels=256 $B\times 256\times H_{in}/4\times H_{in}/4$
+Conv2D, BN, ReLU kernel=3, stride=1, padding=1, $256\to 256$ $B\times 256\times H_{in}/4\times H_{in}/4$
+Self Attention channels=256 $B\times 256\times H_{in}/4\times H_{in}/4$
+Conv2D, BN, ReLU kernel=3, stride=1, padding=1, $256\to 256$ $B\times 256\times H_{in}/4\times H_{in}/4$
+Self Attention channels=256 $B\times 256\times H_{in}/4\times H_{in}/4$
+Conv2D, BN kernel=3, stride=1, padding=1, $256\to 128$ $B\times 128\times H_{in}/4\times H_{in}/4$
+Decoder Block 1 Upsample Target: $H_{in}/2\times H_{in}/2$ $B\times 128\times H_{in}/2\times H_{in}/2$
+Conv2D kernel=3, stride=1, padding=1, $256\to 64$ $B\times 64\times H_{in}/2\times H_{in}/2$
+Decoder Block 2 Upsample Target: $H_{in}\times H_{in}$ $B\times 64\times H_{in}\times H_{in}$
+Conv2D kernel=3, stride=1, padding=1, $128\to C_{out}$ $B\times C_{out}\times H_{in}\times H_{in}$
+Figure 7: Student attention module layout.
 (a) with $A^{i}$
@@
 4 ToyCity Dataset
-The ToyCity dataset addresses the need for a benchmark dataset consisting of real images of multiple objects for Scene AD.
-Unlike MAD, ToyCity consists of (i) substantially more real images and (ii) is not centered around a single object.
+The ToyCity dataset addresses the need for a real image benchmark dataset consisting of multiple objects and multiple views.
+As shown in Table 2, ToyCity is comparable in size to existing prominent AD benchmarks that consist of real images of anomalies.
+Additionally, ToyCity consists of substantially more real images, consists of unconstrained random views, and is not centered around a single object.
 ToyCity features scenes of a toy city with buildings, signs, roads, trees, and miscellaneous items.
-The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure  8.
-Each setup also contains different anomalies, as shown in Figure  8.
-Additionally, each anomaly image has at most one anomaly but each setup has varying anomalies that were added and removed.
-The dataset includes the manually annotated ground truth for anomaly segmentation.
+The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure 9.
+Each setup also contains different anomalies, as shown in Figure 10.
+Additionally, each anomaly image has at most one anomaly, but each setup has varying anomalies that were added and removed as shown in Figure 10.
+Anomalous images contain on average $0.22\%$ anomalous pixels (standard deviation $0.13\%$).
+The percentage of anomalous pixels per image ranges from a minimum of $0.03\%$ to a maximum of $0.78\%$.
 Table 3 provides an overview of the number of images released for ToyCity.
-The dataset’s v1 is the base variant of all four scenes without any augmentations, v2 includes Quer-Aligned Novel Views (QANV) images rendered using Nerfacto, v3 incorporates augmentations using Interpolated Non-Anomalous Views (INV), and v4 combines the data of both v2 and v3, these variations are presented in Table 2.
-ToyCity thus offers a new benchmarking challenge to current state-of-the-art anomaly detection models.
+Each ToyCity has four augmented variations, without any augmentations, with Query-Aligned Novel Views (QANV), with Interpolated Non-Anomalous Views (INV), and combining both QANV and INV; these variations are presented in Table 4.
+ToyCity thus offers 16 total variations that will be evaluated.
+Table 2: Comparison of datasets containing real images of anomalies with ground truth labels for quantitative evaluation.
+Dataset Data Availability Labeled Anomalous Images Multi-View Multi-Object
+AITEX[14] Yes 105 No No
+MPDD[20] Yes 282 No Yes
+MVTec-AD[13] Yes 1258 No No
+MVTec 3D-AD[27] Yes 948 No No
+MVTec LOCO-AD[16] Yes 993 No Yes
+NanoTwice[26] Yes 40 No No
+VisA[15] Yes 1200 No Yes
+MAD-Real[12] Partial 221 Yes No
+ToyCity Yes 1258 Yes Yes
 Figure 9: ToyCity consists of four configurations as shown here.
-We present an example baseline image for each as well as an example of one of several anomalies used.
+We present an example baseline image for each, as well as an example of one of several anomalies used.
 A zoomed comparison is also presented that shows the anomaly highlighted with a red outline.
+Figure 10: Each ToyCity configuration has several anomalies that have been added and removed from the scene. Examples of these anomalies for each ToyCity are shown here by a red bounding box.
 Table 3: Number of non-anomalous and anomalous images in each of the ToyCity dataset scenes, and the number of NVS images generated from each of the proposed methods QANV and IBV
-Name Non-Anomalous Anomalous QANV IBV
-ToyCity 1 316 319 319 3781
-ToyCity 2 341 308 308 3773
-ToyCity 3 307 317 317 3673
-ToyCity 4 314 314 314 3757
+Name Non-Anomalous Anomalous QANV IBV Images
+ToyCity 1 316 319 319 3781 4735
+ToyCity 2 341 308 308 3773 4730
+ToyCity 3 307 317 317 3673 4614
+ToyCity 4 314 314 314 3757 4699
 5 Experiments
 We evaluate OmniAD’s performance, its ablated variants, and several AD methods on ToyCity and MAD to test the localization ability of each method.
-We could only utilize a subset of the MAD dataset that consisted of real images compatible with COLMAP for Structure from Motion (SfM).
+As the camera poses used in MAD-Real [12] for testing were not released publicly, we could only utilize a subset of the MAD dataset that COLMAP could generate camera poses for.
 Each AD method is evaluated quantitatively using the pixel-level $F_{1}$ score and AUROC.
-An ablation study follows to show the significance of the student attention modules on RD.
-Finally a qualitative comparison between OmniAD and RD for anomaly segmentation is presented.
+Additionally, we evaluate the performance of OmniAD on the MVTec-AD dataset [13].
+An ablation study follows to show the significance of the student attention modules on Reverse Distillation.
+Finally, a qualitative comparison between OmniAD and Reverse Distillation for anomaly segmentation is presented.
 5.1 Metrics
-To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$ Score as described by Equation 1.
+To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$,
+$F_{1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}.$ (4)
 The $F_{1}$ score excels in handling class imbalances by merging precision and recall into one comprehensive metric.
 This balance is critical for datasets such as ToyCity and MAD, where anomaly regions typically constitute a small fraction of the total image area.
-Examples of the relatively small anomaly regions within ToyCity can be seen in Figure  8, underscoring the necessity for such a balanced evaluation metric.
+Examples of the relatively small anomaly regions within ToyCity can be seen in Figure  9, underscoring the necessity for such a balanced evaluation metric.
 The imbalance of anomaly regions renders other metrics like accuracy or the area under the receiver operating characteristic curve (AUROC) as misleading.
-$F_{1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}$ (1)
 The reason accuracy is deceptive is that a high accuracy metric can be achieved with a simplistic approach that labels all pixels as non-anomalous.
@@
 However, it does not accurately reflect the model’s ability to precisely localize anomalies, primarily due to the abundance of true negatives.
-Thus we provide both the $F_{1}$ score and AUROC.
-5.2 Implementation Details
-To allow for fair comparison between all anomaly detection methods we use the anomalib [39] library implementation of each anomaly detection method.
-A fixed random seed of 42 was used for all methods.
-All methods were allowed to train for a maximum of 100 epochs and the max achieved $F_{1}$ score was selected as the best model.
-All methods converge well before 100 epochs were completed ensuring fair training.
-Additionally methods were trained with the same batch size.
-Other hyper-parameters were left as default values provided by anomalib for each anomaly detection method.
-5.3 Ablation Studies
-We conducted ablation studies to validate both the OmniAD architecture and the proposed NVS augmentation strategy.
-5.3.1 OmniAD architecture
-OmniAD has two major changes from RD.
-First we change the backbone to ResNeXt and second we introduced branching student attention modules $A^{i}$.
-To test if our changes are significant we test OmniAD with and without ResNeXt (OmniAD-R) and without $A^{i}$ (OmniAD-$A^{i}$) against RD.
-As shown in the $F_{1}$ scores on ToyCity Tables 3 and 4 and MAD Tables 7 we can see a consistent performance increase when $A^{i}$ is used.
-5.3.2 NVS Augmentation
-We test the efficacy of our NVS augmentation strategies by evaluating our metrics with and without each of the augmentations.
-Our hypothesis that generating novel views of the our non-anomalous scene similar to the view point of the query scene will make it easier for OmniAD to identify anomalies is verified through four different variations listed in Table 2.
+Thus, we provide both the $F_{1}$ score and AUROC.
+5.2 Implementation and Training Details
+To ensure a fair comparison and promote reproducibility, all anomaly detection methods evaluated in this study (including CFLOW-AD, FastFlow, CFA, DRÆM, RD, and our proposed OmniAD variants) were implemented using the publicly available anomalib library [37], which has faithfully integrated each method based on original implementations and allows for reproducibility.
+Parameter Settings: For all baseline methods, we utilized the default hyperparameter configurations provided within the anomalib library to match original implementations.
+OmniAD and its ablated versions also inherit these same default hyperparameters from the anomalib implementation of the underlying Reverse Distillation method.
+This approach ensures that performance differences observed are primarily attributable to the architectural modifications (i.e., the ResNeXt backbone and the student attention modules in OmniAD) and the novel view synthesis data augmentation strategies, rather than method-specific hyperparameter tuning.
+Training Procedure: All models were trained for a maximum of 100 epochs, using a batch size of 32 for fair comparison.
+Early stopping was implicitly handled by selecting the model checkpoint that yielded the maximum pixel-level $F_{1}$ score on a validation set, consistent with standard practices.
+We observed that all methods converged well before the 100-epoch limit, ensuring adequate training time.
+Learning rates and model specific training hyperparameters follow the anomalib defaults for each model.
 5.3 Quantitative Evaluation
-In our assessment of diverse AD methods on the ToyCity datasets, we observe a consistent pattern of performance increase OmniAD, OmniAD without ResNext (OmniAD w/o R), and RD when a novel view selection strategy is employed for $F_{1}$ scores on ToyCity and MAD.
-A similar trend can also be seen for AUROC on both ToyCity and MAD Tables 5,6,8.
-Additional experiments on different scenes is provided in the supplemental material.
-The RD with ResNext model and the canonical RD methodology exhibit similar enhancement trends, yet they exhibit smaller gains in comparison to our model.
-The shift from v1 to v2, where anomaly aligned images are introduced, is particularly telling of the significant enhancements in the models’ ability to localize anomalies.
+In our assessment of diverse AD methods for Scene AD using ToyCity and MAD-Real[12], Table 5,6,9, we observe a consistent performance increase for OmniAD compared to Reverse Distillation.
+A similar trend can also be seen for AUROC Table 7,8,10.
+The Reverse Distillation with ResNext model and the canonical Reverse Distillation methodology exhibit similar enhancement trends, yet they exhibit smaller gains in comparison to our model.
+The shift from no augmentation to QANV, where query aligned images are introduced, is particularly telling of the significant enhancements in the models’ ability to localize anomalies due to augmentation.
 Other AD methods such as CFLOW, FastFlow, CFA, and DRÆM show a range of responses to the augmentations, with $F_{1}$ scores experiencing minimal improvements, remaining unchanged, or occasionally declining with the introduction of more complex augmentations.
-This suggests a potential lack of effective utilization of augmented data by these methods in contrast to OmniAD and RD.
-As mentioned metrics such as pixel-level AUROC can often show very high numbers for datasets such as ToyCity and MAD due to the low number of pixels that may be an anomaly.
-This can be seen for both ToyCity Table 5,6 and MAD Table 8.
+This suggests a potential lack of effective utilization of augmented data by these methods in contrast to OmniAD and Reverse Distillation.
+As mentioned, metrics such as pixel-level AUROC can often show very high numbers for datasets, such as ToyCity and MAD, due to the low number of pixels that may be an anomaly.
+This can be seen for both ToyCity Table 7,8 and MAD Table 10.
 Most AD methods will achieve an AUROC very close to 1 while still offering low $F_{1}$ scores.
-Table 2: Four variations are tested for each of the scenes present in ToyCity.
-Variation v1 is un-augmented, v2 is with the Query-Aligned Views, v3 is with the interpolated Non-Anomalous views, and v4 is the combination of Query-Aligned Baseline views and Interpolated Non-Anomalous views.
-Variation Non-Anomalous Anomalous QANV INV
-v1 - -
-v2 -
-v3 -
-v4
+To evaluate the generalizability of OmniAD beyond Scene AD, we also test its performance in a standard fixed-view AD setting using the MVTec-AD dataset[13].
+As indicated in Table 11, OmniAD remains effective in this fixed-view context.
+Although the high AUROC (0.975) achieved by Reverse Distillation for fixed-view AD limits newer methods like OmniAD to incremental improvements until a perfect AUROC is attained.
+OmniAD does demonstrate a more substantial performance gain in terms of the $F_{1}$ score compared to Reverse Distillation and other recent methods such as SimpleNet[51], MambaAD[34], and AnomalyDINO[52].
+Table 4: Four augmentation variations are tested for each of the scenes present in ToyCity.
+Variation Non-Anomalous QANV INV Images
+No Augmentation - - 300+
+QANV Augmentation - 600+
+INV Augmentation - 3600+
+Both QANV and INV 3900+
 5.4 Qualitative Evaluation
-Figure  9 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, RD, which utilizes a ResNeXt backbone.
+Figure 11 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, Reverse Distillation, which utilizes a ResNeXt backbone.
 A threshold was selected for each method that maximizes their pixel-level $F_{1}$ score.
 Three primary failure scenarios are evident.
-First, as depicted in the ToyCity 1 and 4 anomaly instances, RD may erroneously identify arbitrary areas as anomalous.
-Secondly, RD has a propensity to overlook entire anomalous regions, a shortcoming exemplified by the ToyCity 2 anomaly.
-Lastly, RD often overestimates the extent of an anomaly, erroneously marking larger areas than necessary.
-Similar defects in RD can also be seen from the MAD dataset in which it misses an anomaly or overestimates an anomalous region.
-Additional samples from OmniAD for each dataset is presented in Figure 10, showing a heat map of the anomaly score as well as the thresholded segmentation.
-Figure 9: Segmentation performance comparison between OmniAD and RD with ResNeXt backbone across four ToyCity anomalies and 2 objects from MAD.We illustrating three primary failure patterns: misidentifying of non-anomalous areas (ToyCity 1,4m and MAD Pig), overlooking anomalies (ToyCity 2 and Mad Puppy), and overestimation of anomaly areas (ToyCity 3 and MAD Pig).
-Figure 10: Additional segmentations and anomaly scorces from OmniAD.
+First, as depicted in the ToyCity 1 and 4 anomaly instances, Reverse Distillation may erroneously identify arbitrary areas as anomalous.
+Secondly, Reverse Distillation has a propensity to overlook entire anomalous regions, a shortcoming exemplified by the ToyCity 2 anomaly.
+Lastly, Reverse Distillation often overestimates the extent of an anomaly, erroneously marking larger areas than necessary.
+Similar defects in Reverse Distillation can also be seen from the MAD dataset, in which it misses an anomaly or overestimates an anomalous region.
+Additional samples from OmniAD and Reverse Distillation with ResNeXt for each dataset are presented in Figure 12, showing a heat map of the anomaly score as well as the segmentation.
+Additional comparisons between evaluated methods across different anomalies are shown in Figure 13.
+Figure 11: Segmentation performance comparison between the top two evaluated AD methods OmniAD and Reverse Distillation with a ResNeXt backbone across four ToyCity anomalies and objects from MAD.
+We illustrate three primary failure patterns: misidentifying of non-anomalous areas (ToyCity 1,4, and MAD Pig), overlooking anomalies (ToyCity 2 and Mad Puppy), and overestimation of anomaly areas (ToyCity 3 and MAD Pig).
+Other evaluated methods also fail in the same manner as Reverse Distillation with a ResNeXt, but at a higher frequency.
+Figure 12: Additional segmentation and anomaly scores from OmniAD.
+Figure 13: Additional segmentations across all evaluated AD methods.
 Table 5: Comparison of Pixel $F_{1}$ scores on ToyCity 1 and 2
@@
 OmniAD 0.432 0.450 0.497 0.496 0.348 0.599 0.491 0.577
-OmniAD-R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503
-OmniAD-$A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357
-RD 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390
-CFLOW 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082
-FastFlow 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128
-CFA 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081
-DRÆM 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039
+OmniAD w/o R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503
+OmniAD w/o $A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357
+RD [25] 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390
+CFLOW-AD [21] 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082
+FastFlow [22] 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128
+CFA [23] 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081
+DRÆM [24] 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039
 Table 6: Comparison of Pixel $F_{1}$ scores on ToyCity 3 and 4
@@
 OmniAD 0.281 0.384 0.319 0.366 0.379 0.541 0.390 0.523
-OmniAD-R 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
-OmniAD-$A^{i}$ 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
-RD 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
-CFLOW 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
-FastFlow 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
-CFA 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
-DRÆM 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
+OmniAD w/o R 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
+OmniAD w/o $A^{i}$ 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
+RD [25] 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
+CFLOW-AD [21] 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
+FastFlow [22] 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
+CFA [23] 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
+DRÆM [24] 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
 Table 7: Comparison of Pixel AUROC scores on ToyCity 1 and 2
@@
 OmniAD 0.990 0.996 0.999 0.999 0.991 0.998 0.999 0.999
-OmniAD-R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999
-OmniAD-$A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997
-RD 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998
-CFLOW 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868
-FastFlow 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952
-CFA 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984
-DRÆM 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725
+OmniAD w/o R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999
+OmniAD w/o $A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997
+RD [25] 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998
+CFLOW-AD [21] 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868
+FastFlow [22] 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952
+CFA [23] 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984
+DRÆM [24] 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725
 Table 8: Comparison of Pixel AUROC scores on ToyCity 3 and 4
@@
 OmniAD 0.979 0.993 0.997 0.999 0.982 0.993 0.995 0.998
-OmniAD-R 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
-OmniAD-$A^{i}$ 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
-RD 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
-CFLOW 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
-FastFlow 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
-CFA 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
-DRÆM 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
+OmniAD w/o R 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
+OmniAD w/o $A^{i}$ 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
+RD [25] 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
+CFLOW-AD [21] 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
+FastFlow [22] 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
+CFA [23] 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
+DRÆM [24] 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
 Table 9: Comparison of Pixel $F_{1}$ scores on MAD-Real [12] of Puppy, Pig, and Elephant
@@
 OmniAD 0.672 0.714 0.680 0.635 0.334 0.395 0.471 0.524 0.341 0.368 0.374 0.328
-OmniAD-R 0.613 0.677 0.621 0.623 0.302 0.346 0.454 0.441 0.312 0.204 0.297 0.284
-OmniAD-$A^{i}$ 0.634 0.646 0.377 0.363 0.306 0.286 0.343 0.254 0.275 0.298 0.248 0.173
-RD 0.616 0.628 0.520 0.199 0.219 0.203 0.302 0.260 0.262 0.236 0.170 0.192
-CFLOW 0.496 0.466 0.484 0.579 0.162 0.142 0.150 0.147 0.309 0.163 0.214 0.209
-FastFlow 0.646 0.651 0.575 0.598 0.066 0.087 0.043 0.058 0.297 0.285 0.282 0.297
-CFA 0.568 0.576 0.594 0.067 0.294 0.232 0.048 0.029 0.189 0.151 0.117 0.139
-DRÆM 0.042 0.016 0.022 0.018 0.040 0.041 0.072 0.110 0.020 0.041 0.037 0.028
+OmniAD w/o R 0.613 0.677 0.621 0.623 0.302 0.346 0.454 0.441 0.312 0.204 0.297 0.284
+OmniAD w/o $A^{i}$ 0.634 0.646 0.377 0.363 0.306 0.286 0.343 0.254 0.275 0.298 0.248 0.173
+RD [25] 0.616 0.628 0.520 0.199 0.219 0.203 0.302 0.260 0.262 0.236 0.170 0.192
+CFLOW-AD [21] 0.496 0.466 0.484 0.579 0.162 0.142 0.150 0.147 0.309 0.163 0.214 0.209
+FastFlow [22] 0.646 0.651 0.575 0.598 0.066 0.087 0.043 0.058 0.297 0.285 0.282 0.297
+CFA [23] 0.568 0.576 0.594 0.067 0.294 0.232 0.048 0.029 0.189 0.151 0.117 0.139
+DRÆM [24] 0.042 0.016 0.022 0.018 0.040 0.041 0.072 0.110 0.020 0.041 0.037 0.028
 Table 10: Comparison of Pixel AUROC scores on MAD-Real [12] of Puppy, Pig, and Elephant
@@
 OmniAD 0.999 0.999 0.999 0.999 0.988 0.993 0.999 0.999 0.941 0.968 0.973 0.971
-OmniAD-R 0.999 0.993 0.999 0.999 0.988 0.983 0.999 0.999 0.924 0.892 0.933 0.928
-OmniAD-$A^{i}$ 0.999 0.999 0.998 0.999 0.983 0.988 0.997 0.997 0.948 0.946 0.966 0.963
-RD 0.999 0.999 0.999 0.999 0.981 0.982 0.998 0.998 0.936 0.929 0.951 0.956
-CFLOW 0.996 0.996 0.999 0.999 0.962 0.967 0.993 0.990 0.945 0.906 0.947 0.953
-FastFlow 0.999 0.998 0.999 0.999 0.868 0.896 0.965 0.966 0.862 0.914 0.934 0.932
-CFA 0.998 0.998 0.999 0.994 0.977 0.976 0.986 0.979 0.916 0.927 0.934 0.942
-DRÆM 0.819 0.359 0.514 0.524 0.720 0.571 0.684 0.796 0.601 0.496 0.559 0.519
+OmniAD w/o R 0.999 0.993 0.999 0.999 0.988 0.983 0.999 0.999 0.924 0.892 0.933 0.928
+OmniAD w/o $A^{i}$ 0.999 0.999 0.998 0.999 0.983 0.988 0.997 0.997 0.948 0.946 0.966 0.963
+RD [25] 0.999 0.999 0.999 0.999 0.981 0.982 0.998 0.998 0.936 0.929 0.951 0.956
+CFLOW-AD [21] 0.996 0.996 0.999 0.999 0.962 0.967 0.993 0.990 0.945 0.906 0.947 0.953
+FastFlow [22] 0.999 0.998 0.999 0.999 0.868 0.896 0.965 0.966 0.862 0.914 0.934 0.932
+CFA [23] 0.998 0.998 0.999 0.994 0.977 0.976 0.986 0.979 0.916 0.927 0.934 0.942
+DRÆM [24] 0.819 0.359 0.514 0.524 0.720 0.571 0.684 0.796 0.601 0.496 0.559 0.519
+Table 11: Comparison of pixel wise AUROC/$F_{1}$ metrics on the MVTec-AD[13] dataset.
+Method OmniAD RD[25] UniAD[53] SimpleNet[51] AnomalyDINO[52]* MambaAD[34]
+Objects Bottle 0.987/0.755 0.986/0.741 0.983/0.707 0.980/0.727 0.988/0.766
+Cable 0.983/0.633 0.968/0.573 0.974/0.569 0.974/0.600 0.980/0.617
+Capsule 0.984/0.509 0.989/0.499 0.980/0.401 0.989/0.491 0.986/0.475
+Hazelnut 0.993/0.710 0.988/0.608 0.983/0.561 0.979/0.503 0.990/0.639
+Metal Nut 0.987/0.865 0.969/0.795 0.944/0.688 0.988/0.866 0.971/0.806
+Pill 0.983/0.703 0.976/0.667 0.952/0.519 0.985/0.726 0.973/0.660
+Screw 0.993/0.479 0.995/0.472 0.984/0.236 0.993/0.388 0.995/0.517
+Toothbrush 0.992/0.627 0.990/0.609 0.985/0.521 0.985/0.529 0.990/0.607
+Transistor 0.967/0.632 0.906/0.549 0.988/0.585 0.968/0.621 0.968/0.675
+Zipper 0.988/0.686 0.987/0.590 0.965/0.434 0.989/0.654 0.981/0.588
+Textures Carpet 0.986/0.666 0.991/0.609 0.985/0.533 0.980/0.478 0.992/0.634
+Grid 0.994/0.523 0.993/0.477 0.943/0.315 0.988/0.393 0.992/0.488
+Leather 0.995/0.575 0.993/0.467 0.991/0.419 0.992/0.456 0.993/0.495
+Tile 0.956/0.690 0.951/0.597 0.902/0.487 0.972/0.688 0.931/0.523
+Wood 0.955/0.624 0.948/0.495 0.935/0.439 0.940/0.481 0.939/0.477
+Mean 0.983/0.645 0.975/0.583 0.966/0.506 0.980/0.573 0.979/0.618 0.978/0.598
+* Only mean was available from source.
+5.5 Ablation Studies
+We conducted ablation studies to validate both the proposed OmniAD architecture and the effectiveness of the data augmentation strategies.
+5.5.1 OmniAD Architecture Ablation
+To validate the architectural changes proposed in OmniAD and quantify the contribution of each key modification compared to the baseline Reverse Distillation [25], we conducted an ablation study on both OmniAD’s components and data augmentation effects.
+To evaluate the component wise effect of OmniAD we shall refer to the ablation configurations as:
+  • 1.
+    OmniAD: Full model (ResNeXt backbone + $A^{i}$).
+  • 2.
+    OmniAD w/o R: Original Reverse Distillation backbone + $A^{i}$ (isolates $A^{i}$ contribution without ResNeXt).
+  • 3.
+    OmniAD w/o $A^{i}$: ResNeXt backbone without $A^{i}$ (isolates ResNeXt contribution without $A^{i}$).
+  • 4.
+    RD: Baseline Reverse Distillation model (without $A^{i}$ and ResNeXt).
+Results for these configurations are presented in Tables 5, 6, 7, 8, 9, and 10. We aggregate the effects of each component of OmniAD without augmentation in Table 12.
+Table 12: Average Pixel $F_{1}$ Scores across ToyCity and MAD-Real[12] with no augmentation and percentage improvement over baseline RD.
+Method $F_{1}$ Improvement over RD
+OmniAD 0.398 32.5%
+OmniAD w/o R 0.376 25.1%
+OmniAD w/o $A^{i}$ 0.334 11.1%
+RD[25] 0.300 0.0%
+Contribution of Student Attention Modules ($A^{i}$): By comparing configurations with and without $A^{i}$ while keeping the backbone consistent (OmniAD w/o R vs.
+RD), we observe the impact of the student attention module.
+As shown in Table 12, the inclusion of $A^{i}$ leads to a 25.1% gain over RD.
+This supports our hypothesis that expanding the ERF via the student attention modules significantly enhances the model’s ability to localize anomalies.
+Contribution of ResNeXt Backbone: The impact of the ResNeXt backbone is assessed by comparing OmniAD and RD configurations.
+As shown in Table 12, by comparing different backbones (OmniAD vs.
+OmniAD w/o $A^{i}$), using ResNeXt provides an 11.1% gain.
+In summary, the architecture ablation study quantitatively demonstrates that both the ResNeXt backbone and the proposed student attention modules ($A^{i}$) contribute positively to the performance of OmniAD.
+The combination of these components in OmniAD leads to the superior performance observed in our experiments over Reverse Distillation.
+5.5.2 NVS Augmentation Ablation
+To determine the effectiveness of our Novel View Synthesis (NVS) data augmentation techniques, specifically Query-Aligned Non-anomalous Views (QANV) and Interpolated Non-anomalous Views (INV), we performed an ablation study on OmniAD’s performance.
+The underlying assumption was that augmenting the training data with synthesized non-anomalous views, either aligned with query images or interpolated between existing views, would enhance the model’s ability to generalize across different viewpoints and consequently improve anomaly detection.
+Table 13 presents a summary of OmniAD’s average pixel $F_{1}$ scores across all datasets under different augmentation conditions.
+Table 13: Average $F_{1}$ score across ToyCity and MAD-Real[12] for different augmentation strategies.
+Method Augmentation Strategy $F_{1}$ Improvement
+OmniAD Both 0.493 64.33%
+OmniAD QANV 0.493 64.33%
+OmniAD INV 0.460 53.33%
+OmniAD No Aug 0.398 32.67%
+RD[25] No Aug 0.300 0%
+The results in Table 13 clearly indicate that NVS-based augmentation markedly boosts OmniAD’s average performance.
+Both the QANV strategy and the combined use of QANV and INV (Both) achieved the highest average $F_{1}$ score of 0.493, which is approximately a 64.33% relative increase compared to Reverse Distillation without augmentation.
+The INV strategy, when used alone, also demonstrated a significant improvement, yielding an average $F_{1}$ score of 0.460.
+The detailed results (Tables 5, 6, 9) reveal that the best performing augmentation (QANV, INV, or Both) can vary depending on the specific dataset and metric of interest.
+QANV frequently is shown to be the optimal in terms of $F_{1}$ for a specific dataset, while INV and ”Both” are frequently better in terms of AUROC.
+However, in real-world applications of Scene AD, selecting the empirically optimal augmentation strategy per dataset is not feasible due to the absence of ground truth labels.
+Fortunately, combining both QANV and INV offers a robust strategy that does not require dataset-specific optimization.
+The aggregated results across datasets shown in Table 13 shows that the combination of QANV and INV achieves the same high average $F_{1}$ score as QANV across all datasets.
+Thus, while combining QANV and INV may not be the most optimal for a specific dataset, the combination is optimal on average.
 6 Limitations and Future Work
@@
 Note that ToyCity does include partial occlusions, and OmniAD itself does not strictly limit detection to a single anomaly region.
-Second, although ToyCity was captured with real cameras at varying angles and lighting conditions, its tabletop-scale scenes do not fully encompass the complexities of large scale infrastructure.
+Second, although ToyCity was captured with real cameras at varying angles and lighting conditions, its tabletop-scale scenes do not fully encompass the complexities of large-scale infrastructure.
 Furthermore, our data augmentation assumes sufficient coverage for camera localization.
@@
   • 3.
-    Investigate faster 3D reconstruction, localization, and noval view synthesis methods.
-  • 4.
-    Identifying the underlying failure modes in GAN/autoencoder based AD methods under significant viewpoint variation remains an open research question, which we leave for future work.
+    Investigate faster 3D reconstruction, localization, and novel view synthesis methods.
 7 Conclusion
-Our study introduced and addressed a specific anomaly detection (AD) problem we term Scene AD.
-Scene AD deals with pixel-level AD between a non-anomalous scene and an anomalous scene, where a misalignment between non-anomalous and anomalous views is expected, as is often the case in applications in the built environment.
-To challenge and measure the effectiveness of state-of-the-art anomaly detection for Scene AD, we introduced the ToyCity dataset.
-This new benchmark focuses on viewpoint variance in scenes of multiple objects and serves as a testbed for our new method OmniAD.
-We demonstrated through extensive comparisons that OmniAD outperforms other AD methods when applied to the ToyCity dataset as well as the object centric established MAD-Real, establishing a new state-of-the-art for this task.
-Particularly, we found that the incorporation of a student attention module increases the efficacy of reverse distillation for pixel-level AD by expanding the effective receptive filed (ERF).
-Our experiments also demonstrated the benefits of novel view synthesis and camera localization for dataset augmentation in enhancing the robustness and accuracy of OmniAD allowing for an increase in pixel level $F_{1}$ of up to 40% in many scenes.
-We hope that our work will catalyze further advancements, spurring the creation of advanced anomaly detection techniques capable of managing the complexities associated with varying viewpoints in 3D scenes.
+This paper introduced and addressed the task of Scene Anomaly Detection (Scene AD), focusing on localizing anomalies within multi-object scenes captured through unlabeled images exhibiting significant, arbitrary viewpoint variations.
+We identified the limitations of existing datasets and methods, which typically assume fixed cameras or single-object focus.
+To bridge this gap, we presented ToyCity, the first multi-object, multi-view real-image dataset specifically designed to evaluate unsupervised AD methods for Scene AD.
+Furthermore, we demonstrated that established unsupervised AD techniques struggle with pixel-level localization on this challenging dataset.
+We proposed OmniAD, a novel anomaly detection architecture that refines the Reverse Distillation framework.
+By incorporating a more robust ResNeXt backbone and, critically, introducing student attention modules to expand the student decoder’s effective receptive field (ERF), OmniAD significantly enhances sensitivity to localized anomalies across varying perspectives.
+Complementing the architecture, we developed two augmentation strategies, Interpolated Non-anomalous Views (INV) and Query-Aligned Non-anomalous Views (QANV), which utilize a Neural Radiance Field for novel view synthesis (NVS) and hierarchical localization to generate targeted non-anomalous training views.
+Extensive experiments and ablation studies conducted on our ToyCity dataset and the real image subset of MAD validated our approach, showing that OmniAD with QANV and INV improves pixel-level anomaly localization ($F_{1}$ scores) by 64.33% over Reverse Distillation without QANV and INV.
+The results underscored the distinct contributions of the ERF expansion via student attention modules and the effectiveness of NVS-based augmentations.
+We hope our work provides a valuable initial benchmark for future research in view invariant anomaly detection for real-world scenes.
 8 Acknowledgments